### v2

# Competitive Equilibrium for Electricity Markets with Spatially Flexible Loads

Abstract
Electric vehicle charging and geo-distributed datacenters introduce spatially flexible loads (FLs) that couple power, transportation, and datacenter networks. These couplings create a closed-loop feedback between locational marginal prices (LMPs) and decisions of the FL systems, challenging the foundations of conventional competitive equilibrium (CE) in electricity markets. This paper studies a notion of generalized competitive equilibrium (GCE) that aims to capture such price-demand interactions across the interconnected infrastructures. We establish structural conditions under which the GCE preserves key properties of the conventional CE including existence, uniqueness, and efficiency, without requiring detailed knowledge of decision processes for individual FL systems. The framework generalizes to settings where the grid is coupled with multiple FL systems. Stylized examples and case studies on the New York ISO grid coupled with the Sioux Falls transportation and distributed datacenter networks demonstrate the use of our theoretical framework and illustrate the mutual influence among the grid and the studied FL systems.
Index Terms:
Electricity market, Datacenters, Electric Vehicles, Spatial Flexibility.
I Introduction
With the electrification of transportation and the rapid expansion of large-scale, energy-intensive computing infrastructures for AI, power systems are becoming increasingly complex and more tightly integrated with other sectors. In 2023, electric vehicle (EV) electricity consumption in the U.S. surged to 7,596 GWh, nearly five times the level in 2018 [1]. Simultaneously, datacenters accounted for approximately 4% of U.S. electricity demand in 2022, a share projected to reach nearly 6% by 2026 [2]. However, the impact of these new loads extends far beyond this sheer increase in demand. Their nature differs fundamentally from conventional demand, which is fixed at a single location. In contrast, both EVs and geo-distributed datacenters can shift their energy consumption across different locations and this distinctive characteristic defines them as spatially flexible loads (FLs). Furthermore, the collective FLs may depend on external infrastructure networks, such as the transportation network for EVs and compute/communication infrastructure for datacenters, leading to the notion of FL systems that interconnect with the grid.
With the growing prevalence of FLs, understanding their system-wide impacts is essential. In the classical setting with exogenous nodal demand, locational marginal prices (LMPs) at competitive equilibrium (CE) are the benchmark signals aligning private supply/demand choices with social welfare. The existence, uniqueness, and efficiency of the classical CE are well understood. Spatial flexibility complicates this picture: (i) FLs re-site consumption in response to LMPs, and the resulting spatial load profile feeds back into those very LMPs; and (ii) FL systems often comprise many self-interested decision makers coupled through non-electric networks (e.g., transportation, communications). This raises two questions: Can CE be extended to capture this two-way coupling between the grid and FL systems? Under what conditions do existence, uniqueness, and efficiency persist?
I-A Contributions and Paper Organization
To address these questions, we develop a unified framework that couples power-system operation with the spatial load-shifting behavior of FLs such as electrified transportation and geo-distributed datacenters. We first introduce an abstract FL model and define the generalized competitive equilibrium (GCE) for a wholesale electricity market coupled with an FL system (Section II), then analyze its structural properties (Section III). We specialize the framework to EV charging (Section IV) and datacenter loads (Section V), and demonstrate a case where the grid is simultaneously coupled to both FL systems (Section VI). Our contributions are:
  1. i)
    A modeling framework for GCE that unifies spatial flexibility across heterogeneous FL decision models (non-atomic games, finite-player games, and centralized control) and accommodates coupling through external shared infrastructures (e.g., transportation, communication);
  2. ii)
    General conditions for existence, uniqueness, and efficiency of GCE, which translate into concrete, easily verifiable criteria for FL systems for EVs and datacenters;
  3. iii)
    An extension to multiple, concurrent FL systems, providing (to our knowledge) the first rigorous tools to analyze the couplings among the grid and multiple FL systems in a single equilibrium framework.
I-B Related Literature
Existing studies typically model the interaction between the power grid and a specific class of FLs, without accounting for the mutual impact these FLs’ decisions have on each other due to the external infrastructure network that they share. Very few studies treat the coupling between the grid and an FL system (mostly electrified transportation), offering theoretical insights into efficiency or uniqueness properties that are highly problem-dependent. Consequently, the literature lacks a unified theoretical framework that explains the structural coupling and equilibrium properties shared across different FL systems. Market equilibrium with multiple, diverse FL systems connected to the grid simultaneously has not been studied. We survey application-specific literature next.
Studies on power-transportation coupling generally fall into two categories. The first assumes a central coordinator, which minimizes combined transportation and power system costs [3], guides EV routing to alleviate grid congestion [4], or reduces overall operational costs of EVs [5]. The second extends classical transportation models by considering a decentralized setting, where numerous EV owners individually select routes and charging locations to minimize travel costs, thereby influencing grid operation. While some studies omit charging costs in route decisions [6] or treat charging prices as exogenous retail rates [7], Alizadeh et al. [8] model EV charging within a competitive electricity market and capture the bidirectional coupling between LMPs and EV behavior, a setting conceptually closest to ours. This model is also applied in [9, 10] to the distribution power system studies. Building on these models, subsequent research designs charging pricing mechanisms to reduce grid peaks [11], maximize charging station profitability [12, 13], or align user behavior with social welfare [14, 15, 16].
Research on datacenter-power system coupling has largely been unidirectional. The first direction models the grid’s impact on datacenters, focusing on how LMPs influence datacenter workload allocation across geographical locations to minimize electricity costs [17, 18, 19] or carbon emissions [20, 21]. The second direction models the datacenter’s impact on the grid, particularly how spatial flexibility aids demand response [22, 23, 24], or affects market outcomes [25, 26]. While a few recent studies address bidirectional interactions, they generally operate outside the framework of competitive wholesale markets. These studies employ price signals other than LMPs to model the datacenter’s response, including Stackelberg game formulations [27], or auction-based market designs [28] and flexibility pricing schemes [29].
II Unified Framework
For a transmission network with $N$ buses and $M$ lines, we use $i\in\mathcal{N}:=\{1,\cdots,N\}$ to index the buses. The LMP at each bus $i$ is denoted as $\lambda_{i}$, and the vector form of the system-wide LMP is $\bm{\lambda}\in\mathbb{R}^{N}$.
II-A Models of Generators and Stationary Loads
We consider a power system in which each bus is equipped with a generator and a stationary load. The generation at bus $i$ is denoted by $g_{i}$, with feasible set $\mathcal{G}_{i}\subseteq\mathbb{R}$ being non-empty, compact and convex. Let $\mathbf{g}\in\mathbb{R}^{N}$ denote the vector of all generator outputs. Each generator optimizes its output in response to the LMP $\lambda_{i}$ at its bus. Assuming a stictly convex cost function $c_{i}(g_{i})$, the cost-minimization problem for the generator at bus $i$ is given by:
$\min_{g_{i}\in\mathcal{G}_{i}}\ c_{i}(g_{i})-\lambda_{i}g_{i}.$ (1)
We denote the feasible set for the collection of generators by $\mathcal{G}:=\{\mathbf{g}\in\mathbb{R}^{N}:g_{i}\in\mathcal{G}_{i},\ i\in\mathcal{N}\}$, and the total generation cost by $C(\mathbf{g}):=\sum_{i\in\mathcal{N}}c_{i}(g_{i})$.
Meanwhile, each bus also hosts a stationary load with consumption level $\ell_{i}$, which is assumed to be inelastic; let $\bm{\ell}\in\mathbb{R}^{N}$ denote the vector of all such loads.
II-B Model of Spatially Flexible Loads
Unlike stationary loads, FLs can shift their electricity consumption across different geographical locations in response to LMPs. Prime examples include EVs, which choose routes and charging sites, and datacenters, whose users may migrate computational tasks among facilities. These FLs collectively form an FL system, in which decisions of individual FLs impact each other through their impact on shared external infrastructure such as a transportation or datacenter network. We focus on a single FL system coupled with the grid in this section, with the extension to multiple FL systems discussed in Section III-B. Let $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{N_{\mathrm{L}}}$ denote the vector of decisions internal to the FL system, where the feasible set $\mathcal{X}$ is a convex and compact set that encodes the physical and operational limits of the FL system and $N_{\mathrm{L}}$} specifies the dimensionality of the decision space. For simplicity, we refer to $\mathbf{x}$ as FL decisions subsequently.
The FLs’ aggregate impact on the power grid is represented by the vector of net power consumption $\mathbf{s}\in\mathbb{R}^{N}$. To have a general yet tractable model, we make several simplifying assumptions next. Despite appearing to be restrictive, we will demonstrate that the resulting model can encapsulate important FL systems, including EVs and datacenters.
Assumption 1 (Linear Decision-Load Mapping).
The mapping from FL decisions $\mathbf{x}$ to their aggregate grid-level power consumption vector $\mathbf{s}$ is linear, modeled as:
$\mathbf{s}=\mathbf{A}^{\mathrm{FL}}\mathbf{x},$ (2)
for some fixed $\mathbf{A}^{\mathrm{FL}}\in\mathbb{R}^{N\times N_{\mathrm{L}}}$. Denote the set of all feasible aggregate FL consumption vectors by $\mathcal{S}:=\{\mathbf{s}\in\mathbb{R}^{N}\mid\mathbf{s}=\mathbf{A}^{\mathrm{FL}}\mathbf{x},\ \mathbf{x}\in\mathcal{X}\}.$
Furthermore, we stipulate that there exists a function $\Phi(\mathbf{x})$, referred to as the collective preference/disutility function for the FL system, that ranks different FL decisions $\mathbf{x}$ without accounting for the cost of electricity, with a lower $\Phi$ value being better. Notably, the function $\Phi(\mathbf{x})$ may not simply be the sum of the cost/disutility of individual agents in the FL system when multi-agent settings are considered due to their strategic interactions; see Sections IV and V for details. Then, the following is a natural model to characterize the aggregate FL behavior given the LMPs.
Assumption 2 (Convex Best Response Model).
The FL system’s best response is to trade off between its internal preference/disutility with the resulting electricity cost:
$\mathbf{x}^{\star}(\bm{\lambda})=\arg\min_{\mathbf{x}\in\mathcal{X}}\left\{\Phi(\mathbf{x})+\bm{\lambda}^{\top}\mathbf{A}^{\mathrm{FL}}\mathbf{x}\right\}.$ (3)
Further, we assume that $\Phi$ is a strictly convex function over the non-empty, compact, and convex set $\mathcal{X}$.
While the above model defines the FL’s internal choices, a power system operator (SO) only observes the aggregate power consumption, $\mathbf{s}$. Given our grid focus, it is ideal to abstract away internal details of the FL system and create a more succinct model of the FL system that do not explicitly depends on $\mathbf{x}$. To this end, we define $J(\mathbf{s})$ as a value function that maps any given aggregate FL power consumption $\mathbf{s}$ to the minimum FL preference value $\Phi$ achievable given $\mathbf{s}$. Mathematically,
$J(\mathbf{s}):=\min_{\mathbf{x}\in\mathcal{X}}\left\{\Phi(\mathbf{x})\quad\mathrm{s.t.}\quad\mathbf{s}=\mathbf{A}^{\mathrm{FL}}\mathbf{x}\right\}.$ (4)
Under Assumption 2, the value function $J(\mathbf{s})$ is also strictly convex over its domain. As a result, we can formulate the aggregate FL power consumption profile $\mathbf{s}$ as a function of the LMPs $\bm{\lambda}$, denoted by $\bm{\sigma}:\mathbb{R}^{N}\mapsto\mathbb{R}^{N}$, taking the form of
$\bm{\sigma}(\bm{\lambda}):=\arg\min_{\mathbf{s}\in\mathcal{S}}\left\{J(\mathbf{s})+\bm{\lambda}^{\top}\mathbf{s}\right\}.$ (5)
This function $\bm{\sigma}(\bm{\lambda})$ is the key interface between the FL system and the power grid.
II-C Economic Dispatch
The SO solves the following economic dispatch problem to determine both the generation dispatch and the resulting prices, minimizing generator costs:
$\min_{\mathbf{g}\in\mathcal{G},\mathbf{p}\in\mathcal{P}}$ $C(\mathbf{g})$ (6a)
$\mathrm{s.t.}\quad$ $\bm{\lambda}:\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s},$ (6b)
[…]
II-D Classical and Generalized Competitive Equilibrium
Figure 1: Closed-loop feedback between the grid and FL system. See Section III-B for the extension to multiple FL systems.
To formally characterize the market outcome when the FL profile is treated as exogenous, we introduce the following definition of the classical CE.
Definition 1 (Classical CE given FL Profile).
Given aggregate FL power consumption $\mathbf{s}\in\mathbb{R}^{N}$, a pair $(\mathbf{g}^{\star},\bm{\lambda}^{\star})$ constitutes a classical CE if the following conditions are satisfied:
  1. i)
    Individual rationality for generators: For every bus $i\in\mathcal{N}$, the generation decision $g_{i}^{\star}$ solves
    $\min_{g_{i}\in\mathcal{G}_{i}}\;c_{i}(g_{i})-\lambda_{i}^{\star}g_{i}.$
  2. ii)
    Feasibility: The resulting net injection vector $\mathbf{p}:=\mathbf{g}^{\star}-\bm{\ell}-\mathbf{s}$ is feasible, satisfying network constraints $\mathbf{p}\in\mathcal{P}$.
  3. iii)
    Price optimality: The price vector $\bm{\lambda}^{\star}$ is a dual-optimal solution to the economic dispatch problem (6) given $\mathbf{s}$.
To capture the closed-loop feedback between the grid and FL system in Fig. 1, we extend the classical CE as follows.
Definition 2 (Generalized CE).
A triple $(\mathbf{g}^{\star},\mathbf{s}^{\star},\bm{\lambda}^{\star})$ is said to constitute a GCE if
  1. i)
    The pair $(\mathbf{g}^{\star},\bm{\lambda}^{\star})$ constitutes a classical CE given $\mathbf{s}^{\star}$.
  2. ii)
    The FL profile $\mathbf{s}^{\star}$ represents the aggregate best response to the LMPs, i.e., $\mathbf{s}^{\star}\;=\;\bm{\sigma}\bigl(\bm{\lambda}^{\star}\bigr).$
CE in electricity markets is well-studied and is known to exist under standard convexity assumptions. Moreover, it yields a socially optimal allocation and enjoys uniqueness under appropriate conditions [30, 31]. Building on these classical insights, we aim to investigate whether/when these desirable properties extend to the GCE framework, which is formalized and analyzed in the following sections.
III General Results
In this section, we work with general FL systems to establish properties of the GCE. We start by examining the setting with a single FL system in Section III-A and then address the case where multiple FL systems are simultaneously connected to the grid in Section III-B. As we are building on the model in Section II, Assumptions 1 and 2 are in force for all results in this section.
III-A Grid Coupled with Single FL System
We begin by establishing a structural characterization of the GCE, which allows analysis of its existence and uniqueness.
Theorem 1 (Structural Characterization for GCE).
A triple $(\mathbf{g},\mathbf{s},\bm{\lambda})$ is a GCE if and only if it is a primal-dual optimal solution to the following optimization:
$\begin{aligned} \min_{\mathbf{s}\in\mathcal{S},\,\mathbf{g}\in\mathcal{G},\,\mathbf{p}\in\mathcal{P}}\quad&J(\mathbf{s})+C(\mathbf{g})\\ \mathrm{s.t.}\quad\quad\ \ &\bm{\lambda}:\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s}.\end{aligned}$ ($\mathrm{P}_{\mathrm{GCE}}$)
Theorem 1 indicates that the closed-loop feedback between the grid and the FL system can be characterized by a convex optimization problem, and the equilibrium price $\bm{\lambda}^{\star}$ coincides with the Lagrange multiplier of the nodal power-balance constraint in that optimization. Under the scenario when $J(\mathbf{s})$ is available or can be learned from data, ($\mathrm{P}_{\mathrm{GCE}}$) allows us to abstract away individual FL decisions while preserving their aggregate response to LMPs. This formulation captures the collective behavior of a large population of FL through a low-dimensional, system-level representation.
Leveraging Theorem 1, we extablish the existence and uniqueness of the GCE under the stated assumptions.
Corollary 1 (Existence and Uniqueness of GCE).
Define the feasible set for independent decision variables $\mathcal{F}:=\left\{(\mathbf{s},\mathbf{g})\in\mathcal{S}\times\mathcal{G}:\ \exists\,\mathbf{p}\in\mathcal{P}\ \text{s.t.}\ \mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s}\right\}.$ If $\mathcal{F}\neq\emptyset$, a GCE exists. Furthermore, if linear independence constraint qualification holds for ($\mathrm{P}_{\mathrm{GCE}}$), then the GCE is unique.
Given the existence and the uniqueness, we examine the efficiency of the GCE, i.e., whether it supports the social welfare. To that end, we define the following social welfare maximization (SWM) benchmark. In particular, let $\Phi^{\mathrm{SW}}(\mathbf{x})$ denote the sum of individual disutility functions of all agents in the FL system, which may or may not be the same as $\Phi(\mathbf{x})$, especially when agents act selfishly. Then, the SWM can be written as the following optimization problem:
$\min_{\mathbf{x}\in\mathcal{X},\,\mathbf{g}\in\mathcal{G},\,\mathbf{p}\in\mathcal{P}}$ $\Phi^{\mathrm{SW}}(\mathbf{x})+C(\mathbf{g})$ (7)
$\mathrm{s.t.}$ $\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\mathbf{x},$
where we optimize the disutility/cost of both the FL system and the power system in a centralized manner. To facilitate comparison between GCE and SWM, we obtain a characterization of SWM analogous to that of GCE in Theorem 1:
Lemma 1 (Structural Characterization for SWM).
Let $(\mathbf{s}^{\star},\mathbf{g}^{\star},\mathbf{p}^{\star})$ be an optimal solution to:
$\begin{aligned} \min_{\mathbf{s}\in\mathcal{S},\,\mathbf{g}\in\mathcal{G},\,\mathbf{p}\in\mathcal{P}}\quad&J^{\mathrm{SW}}(\mathbf{s})+C(\mathbf{g})\\ \mathrm{s.t.}\quad\quad\ \ \ &\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s},\end{aligned}$ ($\mathrm{P}_{\mathrm{SWM}}$)
where $J^{\mathrm{SW}}(\mathbf{s}):=\min_{\mathbf{x}\in\mathcal{X}}\left\{\Phi^{\mathrm{SW}}(\mathbf{x}):\mathbf{s}=\mathbf{A}^{\mathrm{FL}}\mathbf{x}\right\}.$ Then there exists an $\mathbf{x}^{\star}$ such that ${A}^{\mathrm{FL}}\mathbf{x}^{\star}=\mathbf{s}^{\star}$, where $(\mathbf{x}^{\star},\mathbf{g}^{\star},\mathbf{p}^{\star})$ is an optimal solution to (7).
Comparing ($\mathrm{P}_{\mathrm{GCE}}$) and ($\mathrm{P}_{\mathrm{SWM}}$) in Theorem 1 and Lemma 1, we obtain the following necessary and sufficient condition for the GCE outcome to support social welfare.
Corollary 2 (Efficiency of the GCE).
Let $(\mathbf{s}^{\mathrm{GCE}},\!\mathbf{g}^{\mathrm{GCE}},\!\mathbf{p}^{\mathrm{GCE}})$ be an optimal solution to ($\mathrm{P}_{\mathrm{GCE}}$). Then the GCE is efficient, i.e., $(\mathbf{s}^{\mathrm{GCE}},\!\mathbf{g}^{\mathrm{GCE}})$ solves ($\mathrm{P}_{\mathrm{SWM}}$), if and only if for every $(\mathbf{s},\mathbf{g})\in\mathcal{F}$, there exist subgradients $\bm{\mu}\in\partial J^{\mathrm{SW}}(\mathbf{s}^{\mathrm{GCE}})$ and $\bm{\nu}\in\partial C(\mathbf{g}^{\mathrm{GCE}})$ such that:
$\langle\bm{\mu},\,\mathbf{s}-\mathbf{s}^{\mathrm{GCE}}\rangle+\langle\bm{\nu},\,\mathbf{g}-\mathbf{g}^{\mathrm{GCE}}\rangle\ \geq 0.$ (8)
Corollary 2 establishes the necessary and sufficient condition for ($\mathrm{P}_{\mathrm{GCE}}$) to coincide with the social welfare optimum. A simple sufficient condition is when $J=J^{\mathrm{SW}}$, in which the two optimization problems share the same form. More generally, condition (8) tests, starting at $(\mathbf{s}^{\mathrm{GCE}},\mathbf{g}^{\mathrm{GCE}})$, whether any feasible direction can further improve social welfare. This characterization applies regardless of whether the solution lies in the interior or on the boundary of the feasible set. In the latter case, it is possible that the GCE supports the social welfare even if $J\neq J^{\mathrm{SW}}$; see an example in Section IV-C.
III-B Generalization to Multiple Independent FL Systems
Consider multiple FL systems connected to the same power system, each responding to common LMPs but without direct interaction beyond their shared grid connection, and each satisfying the framework in Section II. Assuming there are $W$ FL systems, we begin by redefining the variables: Let $\mathbf{S}\in\mathbb{R}^{N\times W}$ denote the matrix of electricity consumption, and its $w$-th column $\mathbf{S}_{w}$ represents the load vector of the $w$-th FL system, associated with its own feasible set $\mathcal{S}_{w}$ and value functions $J_{w}(\cdot)$ and $J_{w}^{\mathrm{SW}}(\cdot)$.
Definition 3 (Stacked FL Representation for Multiple Systems).
Consider $W$ FL systems. Define
$\mathbf{s}:=(\mathbf{S}_{1}^{\top},\ldots,\mathbf{S}_{W}^{\top})^{\top},\quad\mathcal{S}:=\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{W},$ (9a)
$J(\mathbf{s}):=\sum_{w=1}^{W}J_{w}(\mathbf{E}_{w}\mathbf{s}),\quad J^{\mathrm{SW}}(\mathbf{s}):=\sum_{w=1}^{W}J_{w}^{\mathrm{SW}}(\mathbf{E}_{w}\mathbf{s}),$ (9b)
where $\mathbf{E}_{w}\in\mathbb{R}^{N\times NW}$ is the block-selection matrix that satisfies $\mathbf{E}_{w}\mathbf{s}=\mathbf{S}_{w}$, i.e., $(\mathbf{E}_{w})_{i,j}=1$ if and only if j=(w1)N+ij=(w-1)N+i, and 0 otherwise.
We have the following property of GCE for this system:
Theorem 2 (GCE with Multiple Independent FL Systems).
Under Definition 3, a triple $(\mathbf{g},\mathbf{s},\bm{\lambda})$ is a GCE of the coupled system if and only if it is primal–dual optimal to ($\mathrm{P}_{\mathrm{GCE}}$) with $J(\mathbf{s})$ and $\mathcal{S}$ as in Definition 3, and with the nodal balance constraint $\mathbf{p}=\mathbf{g}-\bm{\ell}-\sum_{w}\mathbf{E}_{w}\mathbf{s}.$ The equilibrium price $\bm{\lambda}^{\star}$ coincides with the Lagrange multiplier of the nodal balance constraint. Existence and uniqueness of the GCE follow under the same conditions as in Corollary 1. Moreover, the efficiency characterization in Corollary 2 holds with $J(\mathbf{s})$ and $J^{\mathrm{SW}}(\mathbf{s})$ replaced by their separable forms in (9b).
IV Coupled Power and Transportation Networks
In this section, we specialize our general model presented in Section II to an important class of FL systems: electrified transportation systems. The spatial flexibility arises due to the charging location choices of the population of EVs in such systems, which in turn depend on EVs’ route choices given traffic conditions. Section IV-A outlines essential elements to model the electrified transportation system as an FL system, which then enables us to apply our general results to examine the existence, uniqueness, and efficiency of the GCE in the coupled power-transportation systems in Section IV-B. We close the section with an illustrative example of such a system, highlighting a nuanced comparison between GCE and SWM outcomes under various system parameter regions.
IV-A Elements of the FL System for Electrified Transportation
We model the transportation network as a directed graph $\mathcal{G}^{\mathrm{tr}}=(\mathcal{V}^{\mathrm{tr}},\mathcal{E}^{\mathrm{tr}})$, where $\mathcal{V}^{\mathrm{tr}}$ denotes the set of locations (a subset of which host EV charging stations) and $\mathcal{E}^{\mathrm{tr}}$ denotes the set of directed roads connecting these locations. Let $N_{\mathrm{C}}$ denote the number of charging stations. Let there be $K$ origin-destination (OD) pairs, i.e., $\{(o_{1},d_{1}),\ldots,(o_{K},d_{K})\}$ with $(o_{k},d_{k})\in\mathcal{V}^{\mathrm{tr}}\times\mathcal{V}^{\mathrm{tr}}$. For each OD pair $k$, let $\mathcal{R}_{k}^{\mathrm{tr}}$ denote the finite set of feasible routes, each route being an ordered list of edges. Define the set of all routes as $\mathcal{R}^{\mathrm{tr}}=\bigcup_{k=1}^{K}\mathcal{R}_{k}^{\mathrm{tr}}$ and the number of routes as $N_{\mathrm{R}}=|\mathcal{R}^{\mathrm{tr}}|$. Each route is assumed to include exactly one charging station. [^1] If a route contains multiple charging stations, we conceptually split it into distinct routes, each associated with a single charging station along the path. Define the charger-route incidence matrix $\mathbf{A}^{\mathrm{CR}}\in\mathbb{R}^{N_{\mathrm{C}}\times N_{\mathrm{R}}}$, whose entry $(j,r)$ is 1 if the $r$-th route includes the $j$-th charging station, and 0 otherwise. Each column of $\mathbf{A}^{\mathrm{CR}}$ contains exactly one nonzero entry. Similarly, define the charger-bus incidence matrix $\mathbf{A}^{\mathrm{CB}}\in\mathbb{R}^{N_{\mathrm{C}}\times N}$, where entry $(j,i)$ is 1 if the $j$-th charging station is connected to bus $i$. Denote the total number of EVs in the transportation system by $N_{\mathrm{EV}}$. For simplicity, each EV is assumed to charge a fixed amount of energy $q>0$.
IV-A1 Collective decision model
The collective behavior of the electrified transportation system can be modeled via a non-atomic game, where the EV drivers act as the agents. The continuum of these agents is indexed by $\mathcal{J}:=[0,1]$ and scaled to represent a total population of $N_{\mathrm{EV}}$, grouped by their OD pairs. All EVs within the same OD category are treated as identical. For the $k$-th OD pair, the OD demand is denoted by $\rho_{k}>0$, representing the fraction of EVs traveling from $o_{k}$ to $d_{k}$, with the sum satisfying $\sum_{k=1}^{K}\rho_{k}=1$.
IV-A2 FL decisions
For the FL system representing the EV population, the critical internal decisions are route and charging station choices, captured by the traffic flow vector $\mathbf{x}\in\mathbb{R}_{+}^{N_{\mathrm{R}}}$. Here, entry $x_{r}$ denotes the number of EVs selecting the $r$-th route $R_{r}$ and charge at the charging station associated with the route. We define the feasible set of traffic flows as
$\mathcal{X}\!:=\!\big\{\mathbf{x}\in\mathbb{R}_{+}^{N_{\mathrm{R}}}\!\mid\!\sum\nolimits_{R_{r}\in\mathcal{R}_{k}^{\mathrm{tr}}}x_{r}\!=\!\rho_{k}{N_{\mathrm{EV}}},\ \forall k\in\{1,\!\dots,\!K\}\big\}.$ (10)
That is, for each OD pair $k$, the aggregate flow over all feasible routes in $\mathcal{R}_{k}^{\mathrm{tr}}$ must equal the travel demand $\rho_{k}{N_{\mathrm{EV}}}$, and flows are nonnegative on all routes.
Given the traffic flow $\mathbf{x}$}, the aggregate charging load at each power system bus satisfies (2), where matrix $\mathbf{A}^{\mathrm{FL}}$ is given by $\mathbf{A}^{\mathrm{FL}}={q}\cdot(\mathbf{A}^{\mathrm{CB}})^{\top}\mathbf{A}^{\mathrm{CR}}$.
IV-A3 Disutility functions
Each edge $e\in\mathcal{E}^{\mathrm{tr}}$ in the transportation network is associated with a flow-dependent travel cost function $c_{e}:\mathbb{R}^{+}\mapsto\mathbb{R}^{+}$, representing the monetary value of the travel time on edge $e$. The flow on edge $e$, denoted by $y_{e}(\mathbf{x})$, equals the sum of flow from all routes that include $e$:
$y_{e}(\mathbf{x})=\sum\nolimits_{R_{r}\in\mathcal{R}^{\mathrm{tr}}:e\in R_{r}}x_{r},\quad\forall e\in\mathcal{E}^{\mathrm{tr}}.$ (11)
Thus, edge flows are linear functions of the route flow vector $\mathbf{x}$. Since every driver suffers $c_{e}(y_{e}(\mathbf{x}))$ while traveling through edge $e$ with the mass of such drivers being $y_{e}(\mathbf{x})$, the total FL disutility $\Phi^{\mathrm{SW}}(\mathbf{x})$ is then calculated as:
$\Phi^{\mathrm{SW}}(\mathbf{x})=\sum\nolimits_{e\in\mathcal{E}^{\mathrm{tr}}}\!c_{e}\left(y_{e}(\mathbf{x})\right)y_{e}(\mathbf{x}).$ (12)
IV-A4 FL best response model
We are interested in characterizing how the LMPs affect EVs’ route and charging location choices. Given the LMPs, each EV driver selects a route and charging station aiming to minimize their individual cost. The collective outcome of these selfish individual decisions is characterized by the Nash equilibrium of the nonatomic game, commonly referred to as the user equilibrium (UE) in the transportation literature [32]. Specifically, given the LMP $\bm{\lambda}$, the costs associated with a route $R_{r}$ is decomposed into travel and charging components:
Travel cost: $\pi^{\mathrm{travel}}_{r}(\mathbf{x}):=\sum\nolimits_{e\in R_{r}}c_{e}(y_{e}(\mathbf{x})),$ (13a)
Charging cost: $\pi^{\mathrm{charge}}_{r}(\bm{\lambda}):=\bigl((\mathbf{A}^{\mathrm{FL}})^{\top}\bm{\lambda}\bigr)_{r},$ (13b)
Route cost: $\pi_{r}(\mathbf{x},\bm{\lambda})=\pi^{\mathrm{travel}}_{r}(\mathbf{x})+\pi^{\mathrm{charge}}_{r}(\bm{\lambda}),$ (13c)
where the subscript $(\cdot)_{r}$ extracts the vector’s $r$-th entry. Formally, the UE route flow under prices $\bm{\lambda}$, denoted $\mathbf{x}^{\mathrm{UE}}(\bm{\lambda})$, is defined by the condition that for any OD pair $k$, every route $R_{r}\in\mathcal{R}_{k}^{\mathrm{tr}}$ with positive flow $x_{r}^{\mathrm{UE}}>0$ satisfies
$\pi_{r}(\mathbf{x}^{\mathrm{UE}},\bm{\lambda})\leq\pi_{r^{\prime}}(\mathbf{x}^{\mathrm{UE}},\bm{\lambda}),\quad\forall R_{r^{\prime}}\in\mathcal{R}_{k}^{\mathrm{tr}}.$ (14)
In other words, every EV in any used route does not have an unilateral incentive to pick a different route. The aggregate charging load $\mathbf{s}$ induced at the UE is captured by the price-responsive mapping $\mathbf{s}=\bm{\sigma}(\bm{\lambda})=\mathbf{A}^{\mathrm{FL}}\,\mathbf{x}^{\mathrm{UE}}(\bm{\lambda}).$
IV-B Structural Results for Coupled FL and Power System
Although the UE characterization of the FL best response to LMP $\bm{\lambda}$ in (14) does not directly appear in the form of Assumption 2, it can be equivalently expressed as a convex optimization formulation by properly defining $\Phi(\mathbf{x})$:
Lemma 2 (Equivalence of UE and Best-Response Model).
For a fixed vector of LMPs $\bm{\lambda}$, a feasible traffic flow $\mathbf{x}$ is a UE if and only if it solves
$\min_{\mathbf{x}\in\mathcal{X}}\ \Phi(\mathbf{x})+\bm{\lambda}^{\top}\mathbf{A}^{\mathrm{FL}}\mathbf{x},$ (15)
with the FL preference function formulated as follows:
$\Phi(\mathbf{x}):=\sum\nolimits_{e\in\mathcal{E}^{\mathrm{tr}}}\int_{0}^{y_{e}(\mathbf{x})}c_{e}(\xi)\,\mathrm{d}\xi.$ (16)
In particular, $\Phi(\mathbf{x})$ is strictly convex if for every edge $e\in\mathcal{E}^{\mathrm{tr}}$, the travel cost function satisfies $c_{e}(\cdot)\geq 0,$ and $c_{e}(\cdot)$ is continuous and increasing.
Note that the condition in Lemma 2 on edge-level travel cost will probably hold in practice, since the travel time increases with the traffic flow on the edge due to potential traffic congestion. It follows that Assumption 2 holds for electrified transportation systems under the described conditions for $c_{e}(\cdot)$. With $\Phi(\mathbf{x})$ defined in Lemma 2, the associated value function $J(\mathbf{s})$ follows directly from (4), and the structural characterization of GCE in Theorem 1 can be directly applied for coupled power and transportation systems.
IV-B1 Existence and uniqueness of GCE
With $\Phi(\mathbf{x})$ defined in (16), the price-responsive aggregate load mapping $\bm{\sigma}(\bm{\lambda})$ can be equivalently expressed in (5), which coincides with the UE-based response model $\bm{\sigma}(\bm{\lambda})=\mathbf{A}^{\mathrm{FL}}\,\mathbf{x}^{\mathrm{UE}}(\bm{\lambda}).$ By Corollary 1, the GCE exists and is unique. This implies that, once $J(\mathbf{s})$ is available, the GCE of the coupled system can be computed directly from the system-level optimization problem ($\mathrm{P}_{\mathrm{GCE}}$) in Theorem 1, without requiring detailed information on OD demands, routes, and flows. This abstraction substantially reduces the informational burden and computational complexity to characterize the GCE, avoiding the need to work with the high-dimensional FL decision space.
IV-B2 Social welfare property
With $\Phi^{\mathrm{SW}}(\mathbf{x})$ defined in (12), the SWM of the coupled system along with the value function $J^{\mathrm{SW}}(\mathbf{s})$ is the solution to ($\mathrm{P}_{\mathrm{SWM}}$) defined in Lemma 1. With the above characterization, $J(\mathbf{s})$ and $J^{\mathrm{SW}}(\mathbf{s})$ have different forms, which indicates that the GCE generally fails to achieve the social optimum, and the gap between their solutions quantifies the efficiency loss due to selfish user behavior. A detailed comparison between GCE and SWM outcomes is illustrated via an example in Section IV-C.
IV-C Illustrative Example
As depicted in Fig. 2, we consider a two-route transportation network with a common O-D pair so that $\rho_{1}=1$. Each route includes a charging station node, which is connected to a distinct bus in a two-bus power system.
Figure 2: Coupled transportation and 2-bus power system.
The travel cost for each EV user is $\pi^{\mathrm{travel}}_{1}(\mathbf{x})=ax_{1}$ on route 1 and a constant $\pi^{\mathrm{travel}}_{2}(\mathbf{x})=b$ on route 2. Given the LMPs, the total cost of traveling and charging along route 1 is $ax_{1}+\lambda_{1}q$, while the corresponding cost for route 2 is $b+\lambda_{2}q$. The aggregate EV charging loads at buses 1 and 2 are given by $s_{1}=x_{1}q$ and $s_{2}=x_{2}q$, respectively. There are two generators, one at each bus, each with a quadratic cost function: $\gamma_{1}g_{1}^{2}$ for generator 1 and $\gamma_{2}g_{2}^{2}$ for generator 2. There is also a fixed, stationary demand $\ell_{1}$ at Bus 1. The transmission line capacity is $f^{\max}$.
The FL preference function satisfies [^2] Although route 2 has a constant cost, for this simple network we can show the FL preference function $\Phi(\mathbf{x})$ is strictly convex and Assumption 2 holds. $\Phi(x_{1},x_{2})=\tfrac{1}{2}ax_{1}^{2}+bx_{2}$. Following the definition (4), we have $J(s_{1},s_{2})=(a\,s_{1}^{2})/(2q^{2})+(b\,s_{2})/q$ in this example. The optimization-based characterization for the GCE is as follows:
$\min_{\begin{subarray}{c}s_{1},s_{2},g_{1},g_{2},p_{1}\end{subarray}}\$$J(s_{1},s_{2})+\gamma_{1}g_{1}^{2}+\gamma_{2}g_{2}^{2}$ (17a)
$\mathrm{s.t.}\quad\quad$ $\lambda_{1}:p_{1}=g_{1}-\ell_{1}-s_{1},$ (17b)
$\lambda_{2}:-p_{1}=g_{2}-s_{2},$ (17c)
$-{f}^{\max}\leq p_{1}\leq{f}^{\max},$ (17d)
$s_{1}+s_{2}=q{N_{\mathrm{EV}}},$ (17e)
$s_{i}\geq 0,\quad i=1,2,$ (17f)
$0\leq g_{i}\leq g_{i}^{\max},\quad i=1,2.$ (17g)
Meanwhile, the total FL disutility equals $\Phi^{\mathrm{SW}}(x_{1},x_{2})=ax_{1}^{2}+bx_{2}$, and so $J^{\mathrm{SW}}(s_{1},s_{2})=(a\,s_{1}^{2})/(q^{2})+(b\,s_{2})/q$ in this example. Therefore, the welfare maximization problem $(\mathrm{P}_{\mathrm{SWM}})$ takes the following form:
$\min_{\begin{subarray}{c}s_{1},s_{2},g_{1},g_{2},p_{1}\end{subarray}}\$$J^{\mathrm{SW}}(s_{1},s_{2})+\gamma_{1}g_{1}^{2}+\gamma_{2}g_{2}^{2}$ (18a)
[…]
As $J$ and $J^{\mathrm{SW}}$ do not coincide, we investigate whether the GCE supports social welfare by comparing solutions of (17) and (18).
Lemma 3 (Comparison of $(\mathrm{P}_{\mathrm{GCE}})$ and $(\mathrm{P}_{\mathrm{SWM}})$).
For the GCE obtained via solving (17), the following holds.
  1. i)
    There exist parameter values $q,a,b,\gamma_{1},\gamma_{2},f^{\max},g_{1}^{\max}$ and $g_{2}^{\max}$, such that GCE does not support social welfare.
  2. ii)
    Furthermore, if $a>b$, $\gamma_{2}\ll\gamma_{1}$, $g_{2}^{\max}<\gamma_{1}(\ell_{1}+q)/(\gamma_{1}+\gamma_{2})$, and $g_{1}^{\max}>q+\ell_{1}$, then there exist two thresholds, $\bar{f}^{\max}>\underline{f}^{\max}>0$, such that
    1. (a)
      If $f^{\max}>\bar{f}^{\max}$, power system is uncongested in $(\mathrm{P}_{\mathrm{GCE}})$ and $(\mathrm{P}_{\mathrm{SWM}})$, and GCE does not support social welfare.
    2. (b)
      If $\underline{f}^{\max}<f^{\max}\leq\bar{f}^{\max}$, power system congests only in $(\mathrm{P}_{\mathrm{GCE}})$, and GCE does not support social welfare.
    3. (c)
      If $f^{\max}\leq\underline{f}^{\max}$, power system is congested in $(\mathrm{P}_{\mathrm{GCE}})$ and $(\mathrm{P}_{\mathrm{SWM}})$, and GCE support the social welfare.
Lemma 3 shows that, although the GCE generally does not coincide with the social welfare solution, under certain parameter settings, the GCE can nontheless support social welfare even when $J(\mathbf{s})\neq J^{\mathrm{SW}}(\mathbf{s})$. In the parameter regime of Lemma 3-ii), route 1 is less costly than route 2 when its flow is below $b/a$, but becomes more expensive when this threshold is exceeded. Generator 2 is cost-efficient but capacity-limited, while generator 1 is more costly but unconstrained. The lemma further partitions outcomes into three transmission-capacity regimes, each corresponding to a distinct power system congestion pattern: When $f^{\max}$ is sufficiently large, the uncongested system has uniform charging costs, and user choices depend solely on travel costs. In $(\mathrm{P}_{\mathrm{GCE}})$, users selfishly over-utilize the flexible route 1, leading to higher load at bus 1 (with the more expensive generator) compared to the more balanced profile in $(\mathrm{P}_{\mathrm{SWM}})$. In fact, it can be shown that $s_{1}^{\mathrm{GCE}}=2s_{1}^{\mathrm{SWM}}>0$ in Lemma 3-ii)-a). As a result, $(\mathrm{P}_{\mathrm{GCE}})$ relies heavily on the transmission line to import cheaper power from bus 2, making it more prone to line congestion. Thus, for an intermediate range of $f^{\max}$, power system congests only in $(\mathrm{P}_{\mathrm{GCE}})$. Finally, when $f^{\max}$ is small enough so the transmission constraint is binding for both models, the higher LMP at bus 1 discourages drivers from choosing route 1. The charging load at bus 1 can be shown to be identical for both models, satisfying $s_{1}=q-g_{2}^{\max}+f^{\max}$in Lemma 3-ii)-c).
V Coupled Power and Datacenter Networks
This section adapts the general framework from Section II to datacenter networks, another important class of FL systems. The spatial flexibility stems from how computing workloads can be geographically distributed among multiple datacenters. Section V-A introduces the elements to model datacenter operations as an FL system, which then enables us to apply the general results to examine the existence, uniqueness, and efficiency of the GCE in the coupled power-computing systems in Section V-B. The section concludes with a stylized example highlighting how GCE and SWM outcomes diverge under varying system and market parameters.
V-A Elements of the FL System
Suppose there are $N_{\mathrm{D}}$ datacenters in the system, and each datacenter is connected to a power grid bus. We define the datacenter-bus incidence matrix as $\mathbf{A}^{\mathrm{DB}}\in\mathbb{R}^{N_{\mathrm{D}}\times N}$, where entry $(j,i)$ is 1 if the $j$-th charging station is connected to bus $i$ in the power grid. Each unit of computing workload consumes a fixed amount of electricity $q>0$.
V-A1 Collective decision model
We model the setting as a finite-player game, where the agents are $K$ AI companies, denoted by set $\mathcal{K}$, each allocating computing workloads spatially across $N_{\mathrm{D}}$ datacenters. When $K=1$, our model reduces to the case that one company balances workloads across its geographically dispersed datacenters. In this case, the decision model for the FL system is a centralized optimization.
V-A2 FL decisions
The workload allocation strategy of company $k$ is represented by a vector $\mathbf{x}^{(k)}$, where $x_{j}^{(k)}$ denotes the portion of workload assigned to datacenter $j$. Stacking all companies’ strategies yields the FL decision vector $\mathbf{x}\in\mathbb{R}^{KN_{\mathrm{D}}}$: $\mathbf{x}=\bigl(\mathbf{x}^{(1)\top},\dots,\mathbf{x}^{(K)\top}\bigr)^{\top}.$ We define the block summation matrix $\mathbf{A}^{\mathrm{sum}}\in\mathbb{R}^{N_{\mathrm{D}}\times KN_{\mathrm{D}}}$ as $\mathbf{A}^{\mathrm{sum}}=\left(\mathbf{I}_{N_{\mathrm{D}}},\dots,\mathbf{I}_{N_{\mathrm{D}}}\right)$, which aggregates all companies’ allocations to each datacenter via $\mathbf{A}^{\mathrm{sum}}\mathbf{x}$. We denote $\mathcal{X}^{(k)}$ as the feasible set of company $k$’s workload allocation, which characterizes the constraints on computing workload across datacenters, and define $\mathcal{X}:=\prod_{k\in\mathcal{K}}\mathcal{X}^{(k)}$ as the aggregate feasible set for all companies. Hence, the aggregate load profile across datacenters satisfies (2), with the matrix $\mathbf{A}^{\mathrm{FL}}$ given by $\mathbf{A}^{\mathrm{FL}}:=q{(\mathbf{A}^{\mathrm{DB}})}^{\top}\mathbf{A}^{\mathrm{sum}}$.
V-A3 Disutility functions
The utility of company $k$ from allocating workloads is denoted by a function $u_{k}(\mathbf{x}^{(k)};\mathbf{x}^{(-k)})$, where $\mathbf{x}^{(-k)}$ represents the collective strategies of all other companies. This notation captures both the intrinsic value of company $k$’s allocation and the externalities induced by the actions of competitors, e.g., due to increased delays when shared communication/computing infrastructure at the datacenters are congested. The total FL distulity function $\Phi^{\mathrm{SW}}(\mathbf{x})$ is then the negative sum of the utility functions from computing workloads across all AI companies, which satisfies
$\Phi^{\mathrm{SW}}(\mathbf{x})=-\sum\nolimits_{k=1}^{K}u_{k}\big(\mathbf{x}^{(k)};\mathbf{x}^{(-k)}\big).$ (19)
V-A4 FL best response model
We assume that each AI company directly pays for the electricity consumed by its workloads in datacenters, an arrangement consistent with self-operated or colocation datacenters, and serving as a stylized abstraction of electricity bill pass-through in cloud computing environments. The net payoff of company $k$ is then the utility from computing workloads minus the cost of electricity. In response to the LMPs $\bm{\lambda}$, each company solves:
$\max\nolimits_{\mathbf{x}^{(k)}\in\mathcal{X}^{(k)}}\ u_{k}\big(\mathbf{x}^{(k)};\mathbf{x}^{(-k)}\big)-q{\big(\mathbf{A}^{\mathrm{DB}}\bm{\lambda}\big)}^{\top}\mathbf{x}^{(k)}.$ (20)
A Nash Equilibrium (NE) arises when no company can improve its payoff by unilaterally changing its decision. Denote an NE decision as $\mathbf{x}^{\mathrm{NE}}(\bm{\lambda})$. The aggregate price-responsive load $\mathbf{s}$ induced at equilibrium is then characterized by the mapping: $\bm{\sigma}(\bm{\lambda})=\mathbf{A}^{\mathrm{FL}}\,\mathbf{x}^{\mathrm{NE}}(\bm{\lambda}).$
V-B Structural Results for Coupled FL and Power System
To establish the equivalence between the NE characterization of the FL best response to the LMP $\bm{\lambda}$ and Assumption 2, we impose the following conditions under which the preference function $\Phi(\mathbf{x})$ can be defined:
Lemma 4 (Equivalence of NE and Best Response Model).
Suppose each utility function $u_{k}(\mathbf{x}^{(k)};\mathbf{x}^{(-k)})$ is continuously differentiable. Let $\mathbf{F}\!:\!\mathcal{X}\!\to\!\mathbb{R}^{KN_{\mathrm{D}}}$ stack the gradients of each company’s utility $u_{k}$ with respect to their own decision variable $\mathbf{x}^{(k)}$, i.e., $\color[rgb]{0,0,0}\mathbf{F}(\mathbf{x})\!=\!\bigl(\nabla_{\mathbf{x}^{(1)}}u_{1}(\mathbf{x})^{\top},\dots,\nabla_{\mathbf{x}^{(K)}}u_{K}(\mathbf{x})^{\top}\bigr)^{\top}.$} For any fixed LMP vector $\bm{\lambda}$, an NE $\mathbf{x}^{\mathrm{NE}}\in\mathcal{X}$ is then the solution to the variational inequality (VI):
$\left\langle-{\color[rgb]{0,0,0}\mathbf{F}(\mathbf{x})}+q{(\mathbf{A}^{\mathrm{sum}})}^{\top}\mathbf{A}^{\mathrm{DB}}\bm{\lambda},\ \mathbf{x}-\mathbf{x}^{\mathrm{NE}}\right\rangle\geq 0,\ \forall\,\mathbf{x}\in\mathcal{X}.$ (21)
If, in addition, the symmetry condition holds:
$\frac{\partial}{\partial\mathbf{x}^{(\tilde{k})}}\left(\nabla_{\mathbf{x}^{(k)}}u_{k}\right)=\frac{\partial}{\partial\mathbf{x}^{(k)}}\left(\nabla_{\mathbf{x}^{(\tilde{k})}}u_{\tilde{k}}\right),\quad\forall\,k,\tilde{k}\in\mathcal{K},$ (22)
then there exists a function $\Phi(\mathbf{x})$ such that $\nabla_{\mathbf{x}^{(k)}}\Phi(\mathbf{x})=-\nabla_{\mathbf{x}^{(k)}}u_{k}(\mathbf{x}^{(k)};\mathbf{x}^{(-k)})$ for all $k$, and the VI and thus the NE is equivalent to (3) with $\Phi(\mathbf{x})$ defined this way. If each $u_{k}$ is strictly concave and thus $\Phi(\mathbf{x})$ is strictly convex, and each feasible set $\mathcal{X}^{(k)}$ is nonempty, convex, and compact, then the optimization is convex and has a unique solution.
The term $\tfrac{\partial}{\partial\mathbf{x}^{(\tilde{k})}}\!\left(\nabla_{\mathbf{x}^{(k)}}u_{k}\right)$ represent the Jacobian matrix of the gradient of player $k$’s utility with respect to another player’s strategy, representing the cross-partial derivatives $\tfrac{\partial^{2}u_{k}}{\partial\mathbf{x}^{(\tilde{k})}\partial\mathbf{x}^{(k)}}$. The symmetry cross-partial-derivative condition (22) ensures that the game is a potential game [33], permitting the existence of a potential function $\Phi(\mathbf{x})$. This condition frequently holds in datacenter environments, as companies sharing computational resources commonly experience reciprocal performance effects, such as shared queueing delays, leading to symmetric interactions in their utility functions. Concave utility functions $u_{k}$ are also typical in practice, as the marginal value of computing resources generally diminishes with increasing workload. Given the conditions in Lemma 4, Assumption 2 holds for datacenter networks. Given the definition of $\Phi(\mathbf{x})$, the corresponding value function $J(\mathbf{s})$ follows from (4), and the optimization-based characterization for the GCE in Theorem 1 then applies to the setting here. If the potential-game condition (22) is not satisfied, a merit-function VI approach is used; see Appendix D.
V-B1 Existence and uniqueness of GCE
Under the defined preference function $\Phi(\mathbf{x})$, the aggregate price-responsive load mapping $\bm{\sigma}(\bm{\lambda})=\mathbf{A}^{\mathrm{FL}}\,\mathbf{x}^{\mathrm{NE}}(\bm{\lambda})$ can be equivalently characterized by the solution to (5). When the conditions in Lemma 4 are satisfied, the GCE of the coupled power-datacenter system exists and is unique according to Corollary 1. Consequently, once the value function $J(\mathbf{s})$ is specified, the GCE can be directly computed from ($\mathrm{P}_{\mathrm{GCE}}$) in Theorem 1 without the need to track detailed workload allocations at the company level, simplifying the analysis of aggregate system behavior.
V-B2 Social welfare property
Value function $J^{\mathrm{SW}}(\mathbf{s})$ follows (4) with the new form of $\Phi^{\mathrm{SW}}(\mathbf{x})$ in (19) and the SWM of the coupled system is the solution to ($\mathrm{P}_{\mathrm{SWM}}$) defined in Lemma 1. Function $J(\mathbf{s})$ corresponds to decentralized equilibria under self-interested firm behavior, whereas $J^{\mathrm{SW}}(\mathbf{s})$ characterizes centralized coordination that maximizes aggregate utility. An example of their explicit forms and the inefficiencies due to misaligned incentives is detailed in Section V-C.
V-C Illustrative Example
Figure 3: Coupled data center and 2-bus power system.
As illustrated in Fig. 3, the system consists of two AI companies that allocate their computing workloads across two data centers to optimize their utility, i.e., $K=2$. In this setup, the two data centers are connected to power system buses 1 and 2, respectively. We do not enforce capacity limits on the datacenter loads, allowing them to be flexible in their load-shifting strategies. Assume there is no stationary demand at the buses, and the only load is from the datacenters. Consequently, the total demand at bus 1 is $s_{1}=q\big(x_{1}^{(1)}+x_{1}^{(2)}\big)$, and the total demand at bus 2 is $s_{2}=q\big(x_{2}^{(1)}+x_{2}^{(2)}\big)$. Additionally, two generators at bus 1 and bus 2 provide power outputs $g_{1}$ and $g_{2}$, respectively, with quadratic cost function $\gamma_{i}g_{i}^{2}$, $i=1,2$.
In this example, we assume the utility function from computing workload for company $k$, $u_{k}\left(\mathbf{x}^{(k)};\mathbf{x}^{(-k)}\right)$, is quadratic:
$\bm{\eta}_{k}^{\top}\mathbf{x}^{(k)}-\frac{1}{2}{\mathbf{x}^{(k)}}^{\top}\mathbf{M}_{k,k}\mathbf{x}^{(k)}-\big(\mathbf{M}_{k,-k}\mathbf{x}^{(-k)}\big)^{\top}\mathbf{x}^{(k)},$ (23)
where $\bm{\eta}_{k}\in\mathbb{R}^{N_{\mathrm{D}}}$ denotes the baseline utility, reflecting factors such as locational preference and latency penalties associated with geographical distance between datacenters and companies. The symmetric matrix $\mathbf{M}_{k,k}\in\mathbb{R}^{N_{\mathrm{D}}\times N_{\mathrm{D}}}$ captures diminishing returns to scale in resource utilization across datacenters, while $\mathbf{M}_{k,-k}\in\mathbb{R}^{N_{\mathrm{D}}\times(K-1)N_{\mathrm{D}}}$ represents the influence of other companies’ workloads on company $k$’s utility, including queueing delays at each datacenter [34].
Given a vector of LMPs $\bm{\lambda}$, the NE can be expressed as a VI problem that finds a vector $\mathbf{x}^{\star}\in\mathcal{X}$ such that:
$\left\langle\mathbf{M}^{\mathrm{NE}}\mathbf{x}^{\star}-\bm{\eta}+{(\mathbf{A}^{\mathrm{FL}})}^{\top}\bm{\lambda},\ \mathbf{x}-\mathbf{x}^{\star}\right\rangle\geq 0,\ \forall\mathbf{x}\in\mathcal{X},$ (24)
where $\bm{\eta}=(\bm{\eta}_{1}^{\top},\bm{\eta}_{2}^{\top})^{\top}$ stacks the baseline utility vectors, and $\mathbf{M}^{\mathrm{NE}}\in\mathbb{R}^{KN_{\mathrm{D}}\times KN_{\mathrm{D}}}$ captures the inter-company coupling:
$\mathbf{M}^{\mathrm{NE}}=\begin{pmatrix}\mathbf{M}_{1,1}&\mathbf{M}_{1,2}\\ \mathbf{M}_{2,1}&\mathbf{M}_{2,2}\end{pmatrix}$. (25)
Under the symmetric condition (22), $\mathbf{M}^{\mathrm{NE}}$ is required to be symmetric, i.e., $\mathbf{M}^{\mathrm{NE}}=(\mathbf{M}^{\mathrm{NE}})^{\top}$, and the corresponding FL preference function $\Phi(\mathbf{x})$ is given by
$\Phi(\mathbf{x})\!=\frac{1}{2}\mathbf{x}^{\top}\mathbf{M}^{\mathrm{NE}}\mathbf{x}-\bm{\eta}^{\top}\mathbf{x}.$ (26)
[…]
Regarding the SWM benchmark, the detailed characterization of (19) for $\Phi^{\mathrm{SW}}(\mathbf{x})$ in this example is given by:
$\Phi^{\mathrm{SW}}(\mathbf{x})=\frac{1}{2}\mathbf{x}^{\top}\mathbf{M}^{\mathrm{SW}}\mathbf{x}-\bm{\eta}^{\top}\mathbf{x},$ (27)
where matrix $\mathbf{M}^{\mathrm{SW}}\in\mathbb{R}^{KN_{\mathrm{D}}\times KN_{\mathrm{D}}}$ is symmetric positive definite, which satisfies:
$\mathbf{M}^{\mathrm{SW}}:=\mathbf{M}^{\mathrm{NE}}+\left(\mathbf{M}^{\mathrm{NE}}\right)^{\top}-\operatorname{blkdiag}\left(\mathbf{M}_{1,1},\mathbf{M}_{2,2}\right),$ (28)
[…]
The following lemma establishes whether the GCE outcome maximizes social welfare.
Lemma 5 (Characterization of $(\mathrm{P}_{\mathrm{GCE}})$).
Assume that both $\mathbf{M}^{\mathrm{NE}}$ and $\mathbf{M}^{\mathrm{SW}}$ are positive definite. Whether the GCE supports social welfare depends on the cross-company externalities, i.e., the structure of the interaction matrix $\mathbf{M}^{\mathrm{NE}}$:
  1. i)
    If there exists no cross-company externalities, i.e., $\bigl(\mathbf{M}_{1,2}=\mathbf{M}_{2,1}=\mathbf{0}\bigr)$, the GCE outcome coincides with the social-welfare benchmark.
  2. ii)
    For generic cross-terms, i.e., $\bigl(\mathbf{M}_{1,2}\!=\!\mathbf{M}_{2,1}\!\neq\!\mathbf{0}\bigr)$, there exist parameter values $q,\bm{\eta},\mathbf{M}^{\mathrm{NE}},\gamma_{1},\gamma_{2},f^{\max},g_{1}^{\max}$ and $g_{2}^{\max}$ where the GCE does not support social welfare.
  3. iii)
    Suppose grid constraints do not bind, i.e., $f^{\max}=g_{i}^{\max}=\infty$, $i=1,2$, and the AI companies have identical baseline-utility, i.e., $\bm{\eta}_{1}=\bm{\eta}_{2}\propto\mathbf{1}$. If $\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{NE}})^{-1}\mathbf{1}\geq\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{SW}})^{-1}\mathbf{1},$ then LMPs in GCE are higher than those in SWM, i.e., $\lambda_{i}^{\mathrm{GCE}}>\lambda_{i}^{\mathrm{SWM}},$ $i=1,2$.
The lemma shows that in a competitive setting, firms’ self-interested decisions may generate negative externalities, as each firm accounts for shared infrastructure only through its own utility and electricity payments, leading to inefficient spatial allocation of computation. In contrast, a social planner would internalize these externalities and balance workloads across locations to minimize total system cost. In particular, under the condition in Lemma 5-iii), firms tend to induce excessive demand, resulting in higher aggregate demand and consequently higher LMPs.
VI Case Study
We study an integrated system comprising the power grid, transportation network, and datacenter network. This case study illustrates how the framework developed in Section II can be applied to real-world settings, and examines how the resulting GCE captures the interactions and mutual influences among these networks and affects welfare outcomes.
VI-A Configurations of the Power and Multi-FL Systems
The power system configuration is modified from the setting in [27], built upon an 11-zone aggregation of the NYISO network that includes generation data, network parameters, baseline load profiles, and inter-zone distance information. The transportation and datacenter networks are electrically connected to the same power system. The transportation network is adapted from the Sioux Falls road network [35], consisting of 24 nodes, 76 edges. Each transportation node is equipped with an EV charging station that connects electrically to one of the 11 zones in the NYISO-based power system, with each zone containing at least one charging stations; these connections determine $\mathbf{A}^{\mathrm{CB}}$. Five datacenters are placed according to [27] and connected to their zonal power-system buses, deciding $\mathbf{A}^{\mathrm{DB}}$. The GCE is evaluated over discrete one-hour market-clearing periods, with all energy-use activities within each interval assumed to occur simultaneously. The configuration details can be found in Appendix 7.
VI-A1 Electrified transportation system
The transportation system is modified from [35], comprising 182 OD pairs with OD demand $\bm{\rho}$ and hosting a total of 158,080 EVs. For each OD pair, the EV driver chooses between two candidate routes with the shortest free-flow travel times: one assuming charging at the origin and the other assuming charging at the destination, which determines $\mathbf{A}^{\mathrm{CR}}$. The travel cost function on edge $e$ is given by $c_{e}(y_{e}(\mathbf{x}))=c_{e}^{0}\big(1+\kappa_{e}y_{e}(\mathbf{x})\big)$, where $c_{e}^{0}$ represents the base cost, computed as the free-flow travel time multiplied by a value of time of \$17.09/hour [36]. The coefficient $\kappa_{e}$ reflects the marginal effect of traffic volume and is obtained by dividing the delay parameter by the edge capacity in [35]. Each EV requires an identical charging amount $q^{\mathrm{EV}}\in[1,20]$ kWh, corresponding to the energy delivered during one hour of Level 1-2 charging [37].
(a) Cost sensitivity to $q^{\mathrm{EV}}$
with $q^{\mathrm{DC}}=10$ kWh.
(b) Cost sensitivity to $q^{\mathrm{DC}}$ with
$q^{\mathrm{EV}}=20$ kWh.
Figure 4: Impact of FL parameters on generation cost.
VI-A2 Datacenter network
The system includes 11 AI companies, each associated with one of the 11 zones in New York. Each company flexibly allocates its computational workload across the five datacenters, with its utility function following the quadratic form (23). The specifications of parameters in the utility function are documented in Appendix 7. Each standardized workload (e.g., an AI inference/training batch) has an energy requirement of $q^{\mathrm{DC}}\in[0,10]$ kWh. Each AI company optimizes the allocation of its fixed total workload across datacenters to maximize its net payoff, and the total amount of workloads is 364,803.
(a) Internal-system effect
with $q^{\mathrm{DC}}=10$ kWh.
(b) Cross-system impact
with $q^{\mathrm{EV}}=20$ kWh.
Figure 5: Travel cost under increasing EV and datacenter loads.
VI-B GCE-Based Interaction and Welfare Analysis
Using the GCE and SWM formulations developed in Section III, we compute the FL desions, generator dispatch, and LMPs $(\mathbf{g}^{\star},\mathbf{s}^{\star},\bm{\lambda}^{\star})$ under GCE and compare them with the corresponding social welfare optimal outcomes. We then vary key parameters, namely, the transmission line capacities, the EV charging amount $q^{\mathrm{EV}}$ and the energy demand per workload $q^{\mathrm{DC}}$, to examine their impacts on the performance of each subsystem. For comparison, we consider two experimental settings: the congested case, based on the original dataset where limited line capacity causes congestion even without interconnection of the FL systems, and the uncongested case, a synthetic scenario in which all line-capacity constraints are relaxed to ensure an uncongested power system.
Fig. 4 illustrates the impact of FLs’ electricity consumption on total generation cost under GCE and SWM, where network congestion results in higher costs for both models. Under a congested power system, the GCE yields a lower generation cost than the SWM, indicating that self-interested FL behaviors can, counterintuitively, better utilize available generation by exploiting LMP signals. As the FL load increases, the generation cost gap between GCE and SWM widens, reaching 4.30% at the rightmost points of both Fig. 4(a) and Fig. 4(b), where total EV and datacenter loads account for 8.52% and 9.83% of the total load, respectively.
Fig. 5 illustrates how the total travel cost for EVs varies with increasing loads from both EVs and datacenters. In the uncongested power system, travel cost is insensitive to load growth, unlike under congestion. With EV charging demand increasing, the travel cost gap between the GCE and SWM models widens significantly in Fig. 5(a), reflecting increased traffic latency caused by EVs’ self-interested routing. In Fig. 5(b), under the GCE model, rising datacenter load raises the total travel cost of EVs by increasing LMPs and shifting their locational pattern, which redistributes traffic and heightens latency. This reflects a clear negative cross-system externality between the datacenter and transportation networks.
(a) Generation Cost.
(b) Travel Cost.
(c) Computing Utility.
Figure 6: Impact of spatial flexibility with $q^{\mathrm{EV}}=20$ kWh and $q^{\mathrm{DC}}=10$ kWh in the congested power system.
We examine the impact of spatial flexibility of the datacenter network by varying the number of available datacenters from 1 to 5 in Fig. 6 with the total workload from the AI companies fixed, where fewer datacenters indicate less flexibility for AI companies to allocate their workloads. The one- and two-datacenter cases are infeasible because both sites lie within the power system’s most congested zone. In Fig. 6(a), increasing the number of datacenters decreases generation cost, as greater spatial flexibility enables workload redistribution toward more efficient and less congested regions of the grid. Spatial flexibility of the datacenter network likewise reduces EVs’ total travel cost under both GCE and SWM frameworks in Fig. 6(b) by flattening LMP disparities, thus mitigating traffic congestion in the transportation network induced by drivers chasing cheaper charging prices. As shown in Fig. 6(c), computing utility improves with spatial flexibility due to reduced queuing delays. When datacenter locations are restricted, workloads concentrate in a few sites, resulting in negative utility from excessive waiting. Overall, Fig. 6 demonstrates that expanding spatial flexibility benefits all coupled infrastructures even in a self-interested GCE model.
VII Conclusion
This paper introduced a unified framework for electricity markets that explicitly captures the bidirectional interaction between FLs and LMPs through the development and analysis of a GCE. We establish the key structural conditions under which this GCE is guaranteed to exist, be unique, and align with the social welfare optimum. A significant strength of this framework is that these desirable properties hold without requiring the market operator to have explicit knowledge of the FLs’ individual decision-making processes. Through case studies integrating the NYISO grid with the Sioux Falls transportation and datacenter networks, we validate our theoretical findings. These simulations confirm the GCE’s applicability and provide quantitative evidence of the significant mutual influence these coupled systems exert on one another, reinforcing the need for such an integrated equilibrium model.
Appendix A Proofs in Section III
Proof of Theorem 1.
The primal problem ($\mathrm{P}_{\mathrm{GCE}}$) is convex, since $J(\mathbf{s})$ and $C(\mathbf{g})$ are convex and the coupling constraint $\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s}$ is affine. Define the Lagrangian:
$\mathcal{L}^{\mathrm{GCE}}(\mathbf{g},\mathbf{s},\mathbf{p},\bm{\lambda})=J(\mathbf{s})+C(\mathbf{g})+\bm{\lambda}^{\top}(\mathbf{p}-\mathbf{g}+\bm{\ell}+\mathbf{s}).$ (29)
By convex analysis, the KKT conditions are both necessary and sufficient for optimality. Moreover, the KKT system for the primal-dual optimum $(\mathbf{g}^{\star},\mathbf{s}^{\star},\mathbf{p}^{\star},\bm{\lambda}^{\star})$ can be written as the following stationarity and primal feasibility conditions:
$-\partial C(\mathbf{g}^{\star})+\bm{\lambda}^{\star}$ $\in N_{\mathcal{G}}(\mathbf{g}^{\star}),$ (30a)
$-\partial J(\mathbf{s}^{\star})-\bm{\lambda}^{\star}$ $\in N_{\mathcal{S}}(\mathbf{s}^{\star}),$ (30b)
$-\bm{\lambda}^{\star}$ $\in N_{\mathcal{P}}(\mathbf{p}^{\star}),$ (30c)
$\mathbf{0}$ $=\mathbf{p}^{\star}-\mathbf{g}^{\star}+\bm{\ell}+\mathbf{s}^{\star}.$ (30d)
Here, $N_{\mathcal{G}}(\cdot)$, $N_{\mathcal{S}}(\cdot)$, and $N_{\mathcal{P}}(\cdot)$ denote the normal cones to the corresponding feasible sets, where $N_{\mathcal{X}}(x)$ is defined as $N_{\mathcal{X}}(x)=\{\,\nu\mid\nu^{\top}(y-x)\leq 0,\,\forall\,y\in\mathcal{X}\,\}$.
Each condition is equivalent to the optimality condition of a separate subproblem:
[…]
We now show that these KKT conditions are a restatement of the GCE definition.
  1. i)
    Classical CE Condition: A pair $(\mathbf{g}^{\star},\bm{\lambda}^{\star})$ forms a classical CE given $\mathbf{s}^{\star}$ if and only if:
    • Generator individial rationality holds, which is identical to (31a);
    • Feasibility holds, which is identical to (31d);
    • Price optimality holds, the detailed reason is as follows: Given $\mathbf{s}$, the Lagrangian of economic dispatch problem (6) is $\mathcal{L}_{\mathrm{ED}}(\mathbf{g},\mathbf{p};\bm{\lambda})=C(\mathbf{g})+\bm{\lambda}^{\top}\bigl(\mathbf{p}-\mathbf{g}+\bm{\ell}+\mathbf{s}\bigr).$ The dual function decomposes additively:
      $Q(\bm{\lambda})={\inf_{\mathbf{g}\in\mathcal{G}}\bigl\{C(\mathbf{g})-\bm{\lambda}^{\top}\mathbf{g}\bigr\}}+{\inf_{\mathbf{p}\in\mathcal{P}}\bm{\lambda}^{\top}\mathbf{p}}+\bm{\lambda}^{\top}(\bm{\ell}+\mathbf{s}).$
      Hence the dual problem is $\max_{\bm{\lambda}}Q(\bm{\lambda})$. At a dual maximizer $\bm{\lambda}^{\star}$, corresponding primal minimizers is decomposed into two components:
      $\mathbf{g}^{\star}\in\arg\min_{\mathbf{g}\in\mathcal{G}}\{C(\mathbf{g})-\bm{\lambda}^{\star\top}\mathbf{g}\},\ \mathbf{p}^{\star}\in\arg\min_{\mathbf{p}\in\mathcal{P}}\{\bm{\lambda}^{\star\top}\mathbf{p}\},$
      which are exactly (31a) and (31c), respectively.
  2. ii)
    FL Best Response Condition: The FL profile $\mathbf{s}^{\star}$ is the aggregate best response, $\mathbf{s}^{\star}=\bm{\sigma}(\bm{\lambda}^{\star})$, if and only if it solves $\min_{\mathbf{s}\in\mathcal{S}}\{J(\mathbf{s})+\bm{\lambda}^{\star\top}\mathbf{s}\}$, according to the definition in Section II-B. This is identical to (31b).
Since each component of the GCE definition maps bi-directionally to (30), a triple $(\mathbf{g},\mathbf{s},\bm{\lambda})$ is a GCE if and only if it is a primal-dual optimal solution to ($\mathrm{P}_{\mathrm{GCE}}$). ∎
Proof of Corollary 1.
Existence. Since we have:
  1. (C1)
    $\mathcal{G},\mathcal{S}$ and the associated $\mathcal{F}$ are nonempty, convex, and compact;
  2. (C2)
    $J$ and $C$ are continuous, and convex on $\mathcal{S}$ and $\mathcal{G}$, respectively.
Then the objective $J(\mathbf{s})+C(\mathbf{g})$ is continuous on the compact feasible set $\mathcal{F}$. By Weierstrass’ theorem, a minimizer $(\mathbf{g}^{\star},\mathbf{s}^{\star},\mathbf{p}^{\star})$ exists. Since the constraint $\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s}$ is affine, KKT ensures the existence of a multiplier $\bm{\lambda}^{\star}$ satisfying the system (30). Hence a GCE exists.
Uniqueness. Since $J$ is strictly convex on $\mathcal{S}$ and $C$ is strictly convex on $\mathcal{G}$, over $\mathcal{F}$ the objective reduces to the strictly convex function $J(\mathbf{s})+C(\mathbf{g})$ in the variables $(\mathbf{g},\mathbf{s})$ with a convex feasible set $\mathcal{F}$. A strictly convex function has at most one minimizer; hence $(\mathbf{g}^{\star},\mathbf{s}^{\star})$ is unique, and so is $\mathbf{p}^{\star}=\mathbf{g}^{\star}-\bm{\ell}-\mathbf{s}^{\star}$. The uniqueness of the KKT multiplier $\bm{\lambda}^{\star}$ is ensured if linear independence constraint qualification holds for the active constraints at $(\mathbf{g}^{\star},\mathbf{s}^{\star},\mathbf{p}^{\star})$. ∎
Proof of Lemma 1.
Substituting the definition of $J^{\mathrm{SW}}$ into ($\mathrm{P}_{\mathrm{SWM}}$) merges the outer and inner minimizations, yielding exactly the social-welfare program (7) since $\mathbf{s}=\mathbf{A}^{\mathrm{FL}}\mathbf{x}$. Hence, an optimal solution $(\mathbf{s}^{\star},\mathbf{g}^{\star},\mathbf{p}^{\star})$ admits an $\mathbf{x}^{\star}$ attaining $J^{\mathrm{SW}}(\mathbf{s}^{\star})$ with $\mathbf{A}^{\mathrm{FL}}\mathbf{x}^{\star}=\mathbf{s}^{\star}$. Conversely, an $(\mathbf{x}^{\star},\mathbf{g}^{\star},\mathbf{p}^{\star})$ solves (7), establishing the equivalence. ∎
Proof of Corollary 2.
Since $J^{\mathrm{SW}}$ and $C$ are convex and $\mathcal{F}$ is closed and convex, the standard subgradient optimality condition for minimizing a convex function over a convex set (see [38]) implies $(\mathbf{s}^{\star},\mathbf{g}^{\star})$ is optimal for $\min_{(\mathbf{s},\mathbf{g})\in\mathcal{F}}J^{\mathrm{SW}}(\mathbf{s})+C(\mathbf{g})$ if and only if there exist $\bm{\mu}\in\partial J^{\mathrm{SW}}(\mathbf{s}^{\star}),\ \bm{\nu}\in\partial C(\mathbf{g}^{\star})$ such that
$\langle\bm{\mu},\,\mathbf{s}-\mathbf{s}^{\star}\rangle+\langle\bm{\nu},\,\mathbf{g}-\mathbf{g}^{\star}\rangle\geq 0,\ \forall\,(\mathbf{s},\mathbf{g})\in\mathcal{F}.$ (32)
Consequently, $(\mathbf{s}^{\mathrm{GCE}},\mathbf{g}^{\mathrm{GCE}})$ solves ($\mathrm{P}_{\mathrm{SWM}}$) if and only if (8) holds, i.e., the GCE is efficient if and only if (8) holds. ∎
Proof of Theorem 2.
The generalization follows closely from the proof of Theorem 1, with only notational extensions to accommodate multiple FL systems. Under Definition 3, the optimization problem ($\mathrm{P}_{\mathrm{GCE}}$) is convex, since $J(\mathbf{s})=\sum_{w}J_{w}(\mathbf{S}_{w})$ and $C(\mathbf{g})$ are convex and the coupling constraint is affine. The Lagrangian is
$\mathcal{L}^{\mathrm{multi}}(\mathbf{g},\mathbf{s},\mathbf{p},\bm{\lambda})=J(\mathbf{s})+C(\mathbf{g})+\bm{\lambda}^{\top}(\mathbf{p}-\mathbf{g}+\bm{\ell}+\sum_{w}\mathbf{E}_{w}\mathbf{s}).$ (33)
For a primal–dual optimum $(\mathbf{g}^{\star},\mathbf{s}^{\star},\mathbf{p}^{\star},\bm{\lambda}^{\star})$, all KKT conditions are identical to (30), except the $\mathbf{s}$-stationarity (30b), which is:
$\partial J(\mathbf{s}^{\star})-\mathbf{1}\!\otimes\!\bm{\lambda}^{\star}\in N_{\mathcal{S}}(\mathbf{s}^{\star}),$ (34)
where the Kronecker product $\mathbf{1}\!\otimes\!\bm{\lambda}^{\star}\in\mathbb{R}^{WN}$ replicates the same LMP vector $\bm{\lambda}^{\star}$ for each subsystem.
Because $J(\mathbf{s})$ and $\mathcal{S}$ are separable across systems, (34) implies that each subsystem $w$ satisfies
$\mathbf{S}_{w}^{\star}\in\arg\min_{\mathbf{S}_{w}\in\mathcal{S}_{w}}\bigl\{J_{w}(\mathbf{S}_{w})+\bm{\lambda}^{\star\top}\mathbf{S}_{w}\bigr\},$ (35)
which recovers each FL’s best response. The generator and network subproblems remain identical to (31a)-(31c) in the single-system case.
For each element of the GCE definition, we have:
  • FL best response: Each FL subsystem $w$ minimizes its own objective $J_{w}(\mathbf{S}_{w})+\bm{\lambda}^{\top}\mathbf{S}_{w}$ given $\bm{\lambda}$.
  • Classical CE condition: The generator and network decisions $(\mathbf{g},\mathbf{p})$ solve the economic-dispatch and network-feasibility problems given $\bm{\lambda}$; with feasibility constraints replaced by $\mathbf{p}=\mathbf{g}-\bm{\ell}-\sum_{w}\mathbf{E}_{w}\mathbf{s}$ ensures balance.
Hence, $(\mathbf{g}^{\star},\mathbf{s}^{\star},\bm{\lambda}^{\star})$ satisfies the definition of a GCE for multiple independent FL systems.
Because the stacked problem is convex and separable, all arguments for existence, uniqueness, and efficiency carry over verbatim from Corollaries 1 and 2 by substituting $J(\mathbf{s})$ and $J^{\mathrm{SW}}(\mathbf{s})$ with their separable forms in (9b). ∎
Appendix B Proofs in Section IV
Proof of Lemma 2.
Since $c_{e}(\cdot)$ is integrable, define $h_{e}(y):=\int_{0}^{y}c_{e}(\xi)\,\mathrm{d}\xi$. For any $y_{1}\neq y_{2}$ and $\theta\in(0,1)$, let $y_{\theta}:=\theta y_{1}+(1-\theta)y_{2}$. Without loss of generality suppose $y_{1}<y_{2}$, hence $y_{1}<y_{\theta}<y_{2}$. Then if $c_{e}$ is increasing and continuous,
$\theta h_{e}(y_{1})+(1-\theta)h_{e}(y_{2})-h_{e}(y_{\theta})$
$=$ $\theta\!\int_{0}^{y_{1}}\!c_{e}(\xi)\,d\xi+(1-\theta)\!\int_{0}^{y_{2}}\!c_{e}(\xi)\,d\xi-\int_{0}^{y_{\theta}}\!c_{e}(\xi)\,d\xi$
$=$ $-\theta\!\int_{y_{1}}^{y_{\theta}}\!c_{e}(\xi)\,d\xi+(1-\theta)\!\int_{y_{\theta}}^{y_{2}}\!c_{e}(\xi)\,d\xi$
$>$ $-\theta(y_{\theta}-y_{1})\,c_{e}(y_{\theta})+(1-\theta)(y_{2}-y_{\theta})\,c_{e}(y_{\theta})=0,$
where the strict inequality is due to $c_{e}(\xi)<c_{e}(y_{\theta})$ for $\xi\in(y_{1},y_{\theta})$ and $c_{e}(\xi)>c_{e}(y_{\theta})$ for $\xi\in(y_{\theta},y_{2})$. Therefore, $h_{e}(y_{\theta})<\theta h_{e}(y_{1})+(1-\theta)h_{e}(y_{2})$, proving that $h_{e}$ is strictly convex. Since each $y_{e}(\mathbf{x})$ is affine in $\mathbf{x}$ and $\Phi(\mathbf{x})=\sum_{e}h_{e}\big(y_{e}(\mathbf{x})\big)$, it follows that $\Phi(\mathbf{x})$ is continuous and strictly convex in $\mathbf{x}$.
Consider the convex optimization problem (15). Because $\Phi(\mathbf{x})$ is continuous and strictly convex, the KKT condition is necessary and sufficient. Let $\mu_{k}$ denote the multiplier for the equality constraint $\sum_{R_{r}\in\mathcal{R}_{k}^{\mathrm{tr}}}x_{r}=\rho_{k}$, and $\nu_{r}\geq 0$ the multiplier for $x_{r}\geq 0$. Then for each $k$ and $r\in\mathcal{R}_{k}^{\mathrm{tr}}$,
$0\in\partial_{x_{r}}\Phi(\mathbf{x}^{\star})+((\mathbf{A}^{\mathrm{FL}})^{\top}\bm{\lambda})_{r}+\mu_{k}-\nu_{r}.$ (36)
Note that $\partial_{x_{r}}\Phi(\mathbf{x})=\{\pi^{\mathrm{travel}}_{r}(\mathbf{x})\}.$ Since the charging term is linear, $((\mathbf{A}^{\mathrm{FL}})^{\top}\bm{\lambda})_{r}=\pi_{r}^{\mathrm{charge}}(\bm{\lambda})$. Applying complementary slackness $\nu_{r}x_{r}^{\star}=0$ yields two cases:
  • If $x_{r}^{\star}>0$, then $\nu_{r}=0$ and $\pi_{r}^{\mathrm{travel}}(\mathbf{x}^{\star})+\pi_{r}^{\mathrm{charge}}(\bm{\lambda})=-\mu_{k}$, meaning all used routes for OD pair $k$ have equal minimal route cost.
  • If $x_{r}^{\star}=0$, then $\nu_{r}\geq 0$ and $\pi_{r}^{\mathrm{travel}}(\mathbf{x}^{\star})+\pi_{r}^{\mathrm{charge}}(\bm{\lambda})\geq-\mu_{k}$, so unused routes have no lower route cost.
Hence, for each OD pair $k$,
$\pi_{r}(\mathbf{x}^{\star},\bm{\lambda})\leq\pi_{r^{\prime}}(\mathbf{x}^{\star},\bm{\lambda}),\quad\forall R_{r^{\prime}}\in\mathcal{R}_{k}^{\mathrm{tr}},\;\text{with }x_{r}^{\star}>0,$
which is exactly the UE condition (14). Conversely, any $\mathbf{x}^{\star}$ satisfying (14) admits such multipliers $\mu_{k},\nu_{r}$ and therefore satisfies the KKT conditions, proving necessity and sufficiency. This establishes the equivalence between the optimization problem (15) and the UE condition (14). A similar derivation appears in the classical transportation equilibrium theory [32], where the charging component is absent. ∎
Proof of Lemma 3.
Lemma 3-i) can be demonstrated through an example in the parameter regime in Lemma 3-ii).
TABLE I: Illustrative Example: Non-congested Case
$s_{1}^{*}$ $s_{2}^{*}$ $\lambda_{1}^{*}$ $\lambda_{2}^{*}$
GCE $\frac{bq}{a}$ $\frac{(a-b)q}{a}$ $2\gamma_{1}(\ell_{1}+q-g_{2}^{\max})$ $2\gamma_{1}(\ell_{1}+q-g_{2}^{\max})$
SWM $\frac{bq}{2a}$ $\frac{(2a-b)q}{2a}$ $2\gamma_{1}(\ell_{1}+q-g_{2}^{\max})$ $2\gamma_{1}(\ell_{1}+q-g_{2}^{\max})$
Under the parameter regime in Lemma 3-ii), in both ($\mathrm{P}_{\mathrm{GCE}}$) and ($\mathrm{P}_{\mathrm{SWM}}$), the optimal solution for $g_{2}$ satisfies $g_{2}^{*}=g_{2}^{\max}$ and $g_{1}^{*}=\ell_{1}+q-g_{2}^{\max}$. Three mutually exclusive cases for $f^{\max}$ illustrate Lemma 3:
  1. i)
    No congestion $\bigl(f^{\max}\geq g_{2}^{\max}-(a-b)q/a\bigr)$. The line is not congested in both programs. Table I demonstrates the results under this parameter regime. We have $s_{1}^{\mathrm{GCE}}=2s_{1}^{\mathrm{SWM}}$ while the LMPs remain uniform, $\lambda_{1}^{\mathrm{GCE}}=\lambda_{2}^{\mathrm{GCE}}=\lambda_{1}^{\mathrm{SWM}}=\lambda_{2}^{\mathrm{SWM}}$. This result is an example for Lemma 3-i), where $(\mathrm{P}_{\mathrm{GCE}})$ does not support social welfare.
  2. ii)
    Congestion only in GCE $\bigl(g_{2}^{\max}-(2a-b)q/(2a)<f^{\max}<g_{2}^{\max}-(a-b)q/a\bigr)$. The flexible load chosen under the GCE saturates the line, whereas the SWM solution remains uncongested. Hence $s_{1}^{\mathrm{GCE}}=q-g_{2}^{\max}+f^{\max}>s_{1}^{\mathrm{SWM}}$ and the LMPs split,
    $\lambda_{1}^{\mathrm{GCE}}=$ $2\gamma_{1}(\ell_{1}+q-g_{2}^{\max}),$ (37a)
    $\lambda_{2}^{\mathrm{GCE}}=$ $\lambda_{1}^{\mathrm{GCE}}-\frac{a}{q^{2}}{\left(g_{2}^{\max}-\frac{a-b}{a}q-f^{\max}\right)}.$ (37b)
    Therefore, $\lambda_{2}^{\mathrm{GCE}}<\lambda_{1}^{\mathrm{GCE}}=\lambda_{1}^{\mathrm{SWM}}=\lambda_{2}^{\mathrm{SWM}}$.
  3. iii)
    Congestion in both problems $\bigl(f^{\max}\leq g_{2}^{\max}-(2a-b)q/(2a)\bigr)$. The transmission limit binds in both optimisation models, with identical dispatches and identical price profiles, i.e., $s_{1}^{\mathrm{GCE}}=s_{1}^{\mathrm{SWM}}=q-g_{2}^{\max}+f^{\max}$, $\lambda_{1}^{\mathrm{GCE}}=\lambda_{1}^{\mathrm{SWM}}=2\gamma_{1}(\ell_{1}+q-g_{2}^{\max})>\lambda_{2}^{\mathrm{GCE}}=\lambda_{2}^{\mathrm{SWM}}=\lambda_{1}^{\mathrm{GCE}}-{a}/{q^{2}}{\left(g_{2}^{\max}-{(a-b)q}/{a}-f^{\max}\right)}$.
Appendix C Proofs in Section V
Proof of Lemma 4.
Since each utility $u_{k}$ is differentiable and concave and each feasible set $\mathcal{X}^{(k)}$ is convex and compact, the players’ optimization problem in (20) is convex. Its KKT conditions are therefore necessary and sufficient. Collecting the $K$ stationarity conditions yields exactly the VI in (21); Conversely, any $\mathbf{x}^{\star}$ solving (21) satisfies every player’s KKT system and is thus a Nash equilibrium. This one-to-one correspondence follows from Proposition 1.3.4 of [39].
Now suppose the cross-partial symmetry condition (22) holds. By standard results in potential game theory [33], there exists a continuously differentiable scalar function $\Phi(\mathbf{x})$ satisfying $\nabla_{\mathbf{x}^{(k)}}\Phi=-\nabla_{\mathbf{x}^{(k)}}u_{k}$ for all $k$. Substituting this relation into (21) gives
$\big\langle\nabla_{\mathbf{x}}\!\big(\Phi(\mathbf{x}^{\mathrm{NE}})+\bm{\lambda}^{\top}\mathbf{A}^{\mathrm{FL}}\mathbf{x}^{\mathrm{NE}}\big),\,\mathbf{x}-\mathbf{x}^{\mathrm{NE}}\big\rangle\geq 0,\quad\forall\,\mathbf{x}\in\mathcal{X},$
since $\mathbf{A}^{\mathrm{FL}}=q(\mathbf{A}^{\mathrm{DB}})^{\top}\mathbf{A}^{\mathrm{sum}}.$ This is precisely the first-order optimality condition of the convex program $\min_{\mathbf{x}\in\mathcal{X}}\;\Phi(\mathbf{x})+\bm{\lambda}^{\top}\mathbf{A}^{\mathrm{FL}}\mathbf{x},$ proving the equivalence between the VI and the optimization problem (3).
Finally, if each $u_{k}$ is strictly concave and each feasible set $\mathcal{X}^{(k)}$ is nonempty, convex, and compact, then $\mathbf{F}(\mathbf{x})$ is strictly monotone. Thus, the negative mapping gradient mapping $\mathbf{F}(\mathbf{x})=\nabla\Phi(\mathbf{x})$ is strictly monotone, implying that the potential function $\Phi(\mathbf{x})$ is strictly convex. Consequently, the optimization problem is convex. ∎
Proof of Lemma 5.
The only structural difference between optimization frameworks $(\mathrm{P}_{\mathrm{GCE}})$ and $(\mathrm{P}_{\mathrm{SWM}})$ in this example is the interaction matrix, i.e., $\mathbf{M}^{\mathrm{NE}}$ versus $\mathbf{M}^{\mathrm{SW}}$. If there are no cross-company externalities, i.e., $\mathbf{M}_{1,2}=\mathbf{M}_{2,1}=\mathbf{0}$, then $\mathbf{M}^{\mathrm{NE}}=\mathbf{M}^{\mathrm{SW}}$ and hence $J(\mathbf{s})\equiv J^{\mathrm{SW}}(\mathbf{s})$. The two optimization problems coincide, so the outcomes agree, establishing Lemma 5-i).
When there exist cross-company externalities, we write $\mathbf{M}\in\{\mathbf{M}^{\mathrm{NE}},\mathbf{M}^{\mathrm{SW}}\}$ for compactness. Note that for any fixed LMP vector $\bm{\lambda}$, the KKT conditions for the VI problem (24) imply each company’s optimal workload satisfies
$\mathbf{x}^{\star}(\bm{\lambda})=\mathbf{M}^{-1}\!\left(\bm{\eta}-\mathbf{A}^{\mathrm{FL}\top}\bm{\lambda}\right)=\mathbf{M}^{-1}\!\left(\bm{\eta}-q\begin{pmatrix}\bm{\lambda}\\[1.0pt] \bm{\lambda}\end{pmatrix}\!\right)$. (38)
Lemma 5-ii) can be demonstrated through an example in the parameter regime of Lemma 5-iii). Specifically, the generators are assumed to have sufficient capacity, such that their output limits are non-binding. We then assess system equilibrium when the power system is not congested. In this case, the LMPs are same for both buses, i.e., $\lambda^{\star}=\lambda_{1}^{\star}=\lambda_{2}^{\star}=\Gamma(g_{1}^{\star}+g_{2}^{\star})$, where $\Gamma=\frac{2\gamma_{1}\gamma_{2}}{\gamma_{1}+\gamma_{2}}$. The power balance gives $g_{1}^{\star}+g_{2}^{\star}=q\mathbf{1}^{\top}\mathbf{x}^{\star}$. Therefore, substituting (38) leads to the closed-form derivation for LMPs, which is:
$\lambda^{\star}=$ $\frac{\Gamma q\left(\mathbf{1}^{\top}(\mathbf{M})^{-1}\bm{\eta}\right)}{1+\Gamma q^{2}\left(\mathbf{1}^{\top}(\mathbf{M})^{-1}\mathbf{1}\right)}$ (39)
[…]
$\lambda^{\star}=\frac{\Gamma\,q\,\bar{\eta}\,\mathbf{1}^{\top}\mathbf{M}^{-1}\mathbf{1}}{1+\Gamma\,q^{2}\,\bar{\eta}\,\mathbf{1}^{\top}\mathbf{M}^{-1}\mathbf{1}}=\left(q+\frac{1}{\Gamma\,q\,\bar{\eta}\,\mathbf{1}^{\top}\mathbf{M}^{-1}\mathbf{1}}\right)^{-1}.$
Therefore, $\lambda^{\star}$ is increasing in $\mathbf{1}^{\top}\mathbf{M}^{-1}\mathbf{1}$. Therefore, if
$\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{NE}})^{-1}\mathbf{1}\;\geq\;\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{SW}})^{-1}\mathbf{1},$
then
$\lambda_{1}^{\mathrm{GCE}}=\lambda_{2}^{\mathrm{GCE}}\;\geq\;\lambda_{1}^{\mathrm{SWM}}=\lambda_{2}^{\mathrm{SWM}},$
with strict inequality when the above inequality is strict. This proves Lemma 5-iii).
It remains to show that there exist parameters for $\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{NE}})^{-1}\mathbf{1}\;\geq\;\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{SW}})^{-1}\mathbf{1}$ to hold, so the parameter regime discussed above constitutes an example to prove Lemma 5-ii). Particularly, consider the case when the diagonal blocks are the $2\times 2$ identity matrix, i.e., $\mathbf{M}_{1,1}=\mathbf{M}_{2,2}=\mathbf{I}_{2}$. The cross-company interaction blocks are diagonal, $\mathbf{M}_{1,2}=\mathbf{M}_{2,1}=\operatorname{diag}(\vartheta_{1},\vartheta_{2})$, with parameters satisfying $0<\vartheta_{1},\vartheta_{2}<1/2$.
Under the above parameter setting, we can verify that
$\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{NE}})^{-1}\mathbf{1}=$ $2\!\left(\frac{1}{1+\vartheta_{1}}+\frac{1}{1+\vartheta_{2}}\right),$ (40)
$\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{SW}})^{-1}\mathbf{1}=$ $2\!\left(\frac{1}{1+2\vartheta_{1}}+\frac{1}{1+2\vartheta_{2}}\right).$
Hence, condition $\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{NE}})^{-1}\mathbf{1}\geq\mathbf{1}^{\top}(\mathbf{M}^{\mathrm{SW}})^{-1}\mathbf{1}$ is satisfied, leading to higher LMPs under GCE. Since the LMPs differ and $\Gamma(g_{1}^{\star}+g_{2}^{\star})=\lambda^{\star}$, the resulting workload profiles also differ, as $\mathbf{1}^{\top}\mathbf{x}^{\star}=(g_{1}^{\star}+g_{2}^{\star})$ differs. Therefore, $(\mathrm{P}_{\mathrm{GCE}})$ does not support social welfare, which establishes Lemma 5-ii).
Appendix D Non-Potential Game Case for Lemma 4
When condition (22) does not hold, the FL system’s best response to LMP can no longer be written as a convex program in the form of (3). The definitions of $\Phi(\mathbf{x})$ and $J(s)$ in Section II do not apply here. Nevertheless, we can still characterize the price-load mapping and the associated GCE by introducing the merit function of the VI:
Theorem 3 (GCE for an Non-Potential Game).
Suppose each utility function $u_{k}(\mathbf{x}^{(k)};\mathbf{x}^{(-k)})$ is continuously differentiable and let $\mathbf{F}(\mathbf{x})$ follow the definition in Lemma 4. For any fixed LMP vector $\bm{\lambda}\!\in\!\mathbb{R}^{N}$, an NE $\mathbf{x}^{\mathrm{NE}}(\bm{\lambda})$ satisfies (21), which can also be expressed as the minimizer of the following merit-function optimization:
[…]
Assuming $\mathbf{F}$ is strictly monotone on the compact, convex set $\mathcal{X}$, we have that $\mathbf{x}^{\mathrm{NE}}(\bm{\lambda})$ is unique. It follows that the price-load mapping $\mathbf{s}^{\star}=\bm{\sigma}(\bm{\lambda})$ can be expressed as the minimizer of the following optimization:
[…]
$\min_{\mathbf{x}\in\mathcal{X}}\left\{\sup_{\tilde{\mathbf{x}}\in\mathcal{X}}\mathbf{F}(\mathbf{x})^{\top}(\mathbf{x}-\tilde{\mathbf{x}})\ \mathrm{s.t.}\ \mathbf{s}=\mathbf{A}^{\mathrm{FL}}\mathbf{x},\ \tilde{\mathbf{s}}=\mathbf{A}^{\mathrm{FL}}\tilde{\mathbf{x}}\right\}$.
Let $\nabla\mathbf{C}(\mathbf{g})$ denotes the gradient of the total generation cost function $C(\mathbf{g})$ with respect to the generator output vector $\mathbf{g}$. A triple $(\mathbf{g}^{\star},\!\mathbf{s}^{\star},\bm{\lambda}^{\star})$ constitutes a GCE if and only if $(\mathbf{s}^{\star},\!\mathbf{g}^{\star})$ is the minimizer to the following problem
[…]
$\mathrm{s.t.}$ $\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s}\in\mathcal{P},\ \tilde{\mathbf{p}}=\tilde{\mathbf{g}}-\bm{\ell}-\tilde{\mathbf{s}}\in\mathcal{P},$
and the corresponding LMP vector $\bm{\lambda}^{\star}$ is the Lagrange multiplier associated with the nodal constraint in the economic dispatch problem (6) given $\mathbf{s}^{\star}$.
Proof.
Fix an LMP vector $\bm{\lambda}\in\mathbb{R}^{N}$ and define the augmented operator $\mathbf{F}_{\bm{\lambda}}(\mathbf{x})\;:=\;\mathbf{F}(\mathbf{x})\;+\;\bigl(\mathbf{A}^{\mathrm{FL}}\bigr)^{\!\top}\bm{\lambda}.$ The VI problem (21) can be denoted as $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$. Since $\mathbf{F}$ is strictly monotone on the compact, convex set $\mathcal{X}$, $\mathbf{F}_{\bm{\lambda}}$ is strictly monotone as well. Therefore, $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$ has a unique solution according to Theorem 2.3.3 in [39]. The merit function for $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$ is as follows:
[…]
$\Psi_{\bm{\lambda}}$ is non-negative and $\Psi_{\bm{\lambda}}(\mathbf{x})=0$ if and only if $\mathbf{x}$ is the solution to $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}})$; furthermore, strict monotonicity of $\mathbf{F}_{\bm{\lambda}}$ implies $\Psi_{\bm{\lambda}}$ is strictly convex. Consequently, its unique minimizer is $\mathbf{x}^{\mathrm{NE}}(\bm{\lambda})$, resulting in (41). The optimal FL load profile $\mathbf{s}^{\star}$ is then the minimizer of the following problem:
[…]
$\mathrm{s.t.}\quad$ $\mathbf{A}^{\mathrm{FL}}\mathbf{x}=\mathbf{s},\quad\mathbf{A}^{\mathrm{FL}}\tilde{\mathbf{x}}=\tilde{\mathbf{s}}$. (44b)
[…]
Denote $\mathbf{z}=\big(\mathbf{x}^{\top},\,\mathbf{g}^{\top}\big)^{\top}$, and define the block operator
$\mathbf{F}_{\mathbf{C}}(\mathbf{z}):=\begin{bmatrix}\mathbf{F}(\mathbf{x})\\ \nabla C(\mathbf{g})\end{bmatrix}$,
and the convex feasible set
$\Omega\;:=\;\bigl\{\mathbf{z}\mid\mathbf{z}\in\mathcal{Z}=\mathcal{X}\times\mathcal{G},\ \mathbf{p}(\mathbf{z})=\mathbf{g}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\mathbf{x}\!\in\!\mathcal{P}\bigr\}.$
Next, we illustrate the relationship between the characterization of GCE and the solution of $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$. Since $C$ is convex and $\mathbf{F}$ is monotone, $\mathbf{F}_{\mathbf{C}}$ is a monotone operator on the convex set $\Omega$. By Proposition 1.3.4 of [39], a pair $(\mathbf{x}^{\star},\mathbf{g}^{\star})$ solves $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$ if and only if there exists a multiplier $\bm{\lambda}^{\star}$ associated with the coupling constraints $\mathbf{g}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\mathbf{x}\!\in\!\mathcal{P}$ such that the triplet $(\mathbf{x}^{\star},\mathbf{g}^{\star},\bm{\lambda}^{\star})$ satisfies the KKT conditions for the VI, which consist of three parts:
  1. i)
    Stationarity: $\mathbf{0}\in\mathbf{F}_{\mathbf{C}}(\mathbf{z}^{\star})-\nabla\mathbf{p}(\mathbf{z}^{\star})^{\top}\bm{\lambda}^{\star}+N_{\mathcal{Z}}(\mathbf{z}^{\star}).$
  2. ii)
    Primal Feasibility: $\mathbf{p}(\mathbf{z}^{\star})\in\mathcal{P}.$
  3. iii)
    Complementarity: $\bm{\lambda}^{\star}\in N_{\mathcal{P}}(\mathbf{p}(\mathbf{z}^{\star})).$
These conditions map directly to the GCE definition:
  1. i)
    FL system’s best response: The pair $(\mathbf{x}^{\star},\bm{\lambda}^{\star})$ satisfies the KKT conditions for the FL’s NE, expressed as $\operatorname{VI}(\mathcal{X},\mathbf{F}_{\bm{\lambda}^{\star}})$. This implies $\mathbf{x}^{\star}$ is the optimal response to the price $\bm{\lambda}^{\star}$, and the resulting load profile $\mathbf{s}^{\star}:=\mathbf{A}^{\mathrm{FL}}\mathbf{x}^{\star}$ satisfies the price-load mapping $\bm{\sigma}(\bm{\lambda}^{\star})=\mathbf{s}^{\star}$.
  2. ii)
    Classical CE conditions:
    • The pair $(\mathbf{g}^{\star},\bm{\lambda}^{\star})$ satisfies the KKT conditions for generators’ individual rationality in (1).
    • Feasibility holds as $\mathbf{p}^{\star}=\mathbf{g}^{\star}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\mathbf{x}^{\star}\in\mathcal{P}$.
    • Price optimality holds as, $\bm{\lambda}^{\star}\in N_{\mathcal{P}}(\mathbf{p}^{\star})$.
The above conditions, which are the necessary and sufficient KKT conditions for a solution to $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$, are precisely the defining criteria for a GCE. The tuple $(\mathbf{g}^{\star},\mathbf{s}^{\star},\bm{\lambda}^{\star})$ derived from the VI solution $(\mathbf{x}^{\star},\mathbf{g}^{\star},\bm{\lambda}^{\star})$ thus constitutes a GCE. Conversely, any GCE $(\mathbf{g}^{\star},\mathbf{s}^{\star},\bm{\lambda}^{\star})$, along with the corresponding $\mathbf{x}^{\star}$ satisfying $\mathbf{s}^{\star}=\mathbf{A}^{\mathrm{FL}}\mathbf{x}^{\star}$, fulfills all KKT conditions above for VI. Therefore, the correspondence between the solutions of $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$ and the set of GCEs is one-to-one.
Finally, we express $\operatorname{VI}(\Omega,\mathbf{F}_{\mathbf{C}})$ into an optimization problem:
[…]
$\mathrm{s.t.}\quad$ $\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\mathbf{x}\in\mathcal{P},$ (45b)
$\tilde{\mathbf{p}}=\tilde{\mathbf{g}}-\bm{\ell}-\mathbf{A}^{\mathrm{FL}}\tilde{\mathbf{x}}\in\mathcal{P},$ (45c)
[…]
$\mathrm{s.t.}\quad$ $\mathbf{s}=\mathbf{A}^{\mathrm{FL}}\mathbf{x},$ (46b)
$\tilde{\mathbf{s}}=\mathbf{A}^{\mathrm{FL}}\tilde{\mathbf{x}}$, (46c)
$\mathbf{p}=\mathbf{g}-\bm{\ell}-\mathbf{s}\in\mathcal{P},$ (46d)
$\tilde{\mathbf{p}}=\tilde{\mathbf{g}}-\bm{\ell}-\tilde{\mathbf{s}}\in\mathcal{P}.$ (46e)
[…]
Appendix E Case Study Configuration Details
Figure 7: The interconnected (a) power system, (b) transportation system, and (c) datacenter networks.
The 11 zones of the power system, indexed by A-K, are shown in Fig. 7(a). Each zone represents a major operating region (e.g., WEST, CENTRL, N.Y.C), and the transmission lines connecting them indicate the network topology of the regional grid. The coefficient of the quadratic term in the generation cost is set to be ten times larger than that in [27], reflecting a steeper marginal generation cost profile.
The transportation system, shown in Fig. 7(b), consists of 24 nodes representing charging stations as well as the origins and destinations of EV trips. These nodes are grouped into the same 11 regions (A-K) as the power system, with shaded blue areas denoting their geographic correspondence and coupling to the power zones. Arrowed lines indicate the bidirectional links between nodes, representing the road connections that form the transportation network. To better capture real-world congestion effects, the travel demand of each OD pair is adjusted, the free-flow travel time is tripled, and the delay parameter $\kappa_{e}$ is amplified by a factor of ten. The adjustment of the generator quadratic cost coefficients and delay parameters is introduced to increase the sensitivity of marginal generation cost/per driver travel cost to the relocation of traffic flow.
In Fig. 7(c), the geographic locations of the 11 AI companies and 5 datacenters are marked on the map. The solid lines link each AI company and datacenter to their respective regional indices (A-K), indicating their regional associations. Specifically, five datacenters are located in the WEST, CENTRL, NORTH, MILLWD, and N.Y.C regions, consuming electricity locally within their respective zones. For the datacenter network, the dashed lines shown for one representative company illustrate that each AI company can connect to and dispatch workloads across all five datacenters, with different colors corresponding to different datacenters.
TABLE II: Parameters in AI Companies’ Utility Functions (23).
Param Description Value
$\tau_{j}^{(k)}$ Company $k$’s willingness to pay for datacenter $j$ $10^{1}\cdot\mathcal{U}(1,2)$
$\gamma_{k}$ Company $k$’s sensitivity to distance $10^{-1}\cdot\mathcal{U}(2,4)$
$d_{j}^{(k)}$ Distance from company $k$ to datacenter $j$ based on [27]
$\omega_{k}$ Diminishing returns coefficient $10^{-4}\cdot\mathcal{U}(1,5)$
$\alpha_{j}$ Queueing coefficient of datacenter $j$ $10^{-4}\cdot\mathcal{U}(1,2)$
$\zeta_{k}$ Total workload of user $k$ $10^{4}\cdot\mathcal{U}(1,5)$
In (23), each component $\eta_{k,j}$ of the baseline utility vector represents company $k$’s willingness to pay (or preference) for datacenter $j$, denoted by $\tau^{(k)}_{j}$, minus a communication cost proportional to the distance $\gamma_{k}d^{(k)}_{j}$ between them. The quadratic cost terms arise from datacenter queuing effects [34]. At each datacenter $j$, a per-job waiting cost increases proportionally with the total load, $\sum_{\tilde{k}}x_{j}^{(\tilde{k})}$, governed by a queuing coefficient $\alpha_{j}\geq 0$. The total waiting cost for company $k$ is then given by $\sum_{j}\alpha_{j}x_{j}^{(k)}\sum_{\tilde{k}}x_{j}^{(\tilde{k})}$, establishing the interaction between the workloads of competing companies. This interaction directly defines the off-diagonal blocks as $\mathbf{M}_{k,\tilde{k}}=\operatorname{diag}(\alpha_{1},\ldots,\alpha_{N^{\mathrm{D}}})$ for every $\tilde{k}\neq k$, and the row block matrix $\mathbf{M}_{k,-k}$ horizontally concatenates all the individual interaction matrices $\mathbf{M}_{k,\tilde{k}}$. This queuing cost is then combined with an internal cost that captures diminishing marginal returns, modeled with parameter $\omega_{k}>0$ in the quadratic form $\frac{1}{2}\omega_{k}\sum_{j}(x_{j}^{(k)})^{2}$. Then the diagonal blocks satisfy $\mathbf{M}_{k,k}=\operatorname{diag}(2\alpha_{1}+\omega_{k},\ldots,2\alpha_{N^{\mathrm{D}}}+\omega_{k})$. This formulation guarantees the utility function’s concavity, as the requirement $\mathbf{M}_{k,k}\succeq 0$ is always satisfied. The parameters of the utility functions in our simulations are selected according to the aforementioned criteria and summarized in Table II, where $\mathcal{U}(\cdot)$ denotes the uniform distribution. Once the random components of the parameters in the table are determined, these parameters remain fixed throughout all experiments for consistency.
When we vary the number of datacenters from 1 to 5 for assessing the impact of spatial flexibility, we connect each AI company to an expanding set of datacenters, starting from the datacenter in Zone J, and then followed by that in Zones H, D, C, and A, with the order described. This ordering reflects a progressive expansion of spatial flexibility, where each newly added datacenter provides additional siting options for AI companies to allocate their workloads.